LambdaMOO Programmer’s Manual

For LambdaMOO Version 1.8.0p6 March 1997

by Pavel Curtis aka Haakon aka Lambda

Copyright \begin{matrix}c \\ ◯ \end{matrix} 1991, 1992, 1993, 1995, 1996 by Pavel Curtis. Copies of the electronic source for this document can be obtained using anonymous FTP on the In-

ternet. At the site ftp.research.att.com the files are dist/eostrom/MOO/ProgrammersManual.*;

several different file formats are provided, including HTML, Texinfo, plain text, and PostScript. Permission is granted to make and distribute verbatim copies of this manual provided the copyright notice and this permission notice are preserved on all copies. Permission is granted to copy and distribute modified versions of this manual under the conditions for verbatim copying, provided that the entire resulting derived work is distributed under the terms of a permission notice identical to this one. Permission is granted to copy and distribute translations of this manual into another language, under the above conditions for modified versions, except that this permission notice may be stated in a translation approved by the author.

1 Introduction

LambdaMOO is a network-accessible, multi-user, programmable, interactive system well-suited to the construction of text-based adventure games, conferencing systems, and other collaborative software. Its most common use, however, is as a multi-participant, low-bandwidth virtual reality, and it is with this focus in mind that I describe it here.

Participants (usually referred to as players) connect to LambdaMOO using Telnet or some other, more specialized, client program. Upon connection, they are usually presented with a welcome message explaining how to either create a new character or connect to an existing one. Characters are the embodiment of players in the virtual reality that is LambdaMOO.

Having connected to a character, players then give one-line commands that are parsed and interpreted by LambdaMOO as appropriate. Such commands may cause changes in the virtual reality, such as the location of a character, or may simply report on the current state of that reality, such as the appearance of some object.

The job of interpreting those commands is shared between the two major components in the LambdaMOO system: the server and the database. The server is a program, written in a standard programming language, that manages the network connections, maintains queues of commands and other tasks to be executed, controls all access to the database, and executes other programs written in the MOO programming language. The database contains representations of all the objects in the virtual reality, including the MOO programs that the server executes to give those objects their specific behaviors.

Almost every command is parsed by the server into a call on a MOO procedure, or verb, that actually does the work. Thus, programming in the MOO language is a central part of making non-trivial extensions to the database and thus, the virtual reality.

In the next chapter, I describe the structure and contents of a LambdaMOO database. The following chapter gives a complete description of how the server performs its primary duty: parsing the commands typed by players. Next, I describe the complete syntax and semantics of the MOO programming language. Finally, I describe all of the database conventions assumed by the server.

Note: This manual describes only those aspects of LambdaMOO that are entirely independent of the contents of the database. It does not describe, for example, the commands or programming interfaces present in the LambdaCore database.

2 The LambdaMOO Database

In this chapter, I begin by describing in detail the various kinds of data that can appear in a LambdaMOO database and that, therefore, MOO programs can manipulate. In a few places, I refer to the LambdaCore database. This is one particular LambdaMOO database, created every so often by extracting the “core” of the current database for the original LambdaMOO.

Note: The original LambdaMOO resides on the host lambda.parc.xerox.com (the numeric address for which is 192.216.54.2), on port 8888. Feel free to drop by! A copy of the most recent release of the LambdaCore database can be obtained by anonymous FTP from host ftp.parc.xerox.com in the directory pub/MOO.

2.1 MOO Value Types

There are only a few kinds of values that MOO programs can manipulate:

• integers (in a specific, large range)

• real numbers (represented with floating-point numbers)

• strings (of characters)

• objects (in the virtual reality)

• errors (arising during program execution)

• lists (of all of the above, including lists)

MOO supports the integers from −2^31 (that is, negative two to the power of 31) up to 2^31 − 1 (one less than two to the power of 31); that’s from −2147483648 to 2147483647, enough for most purposes. In MOO programs, integers are written just as you see them here, an optional minus sign followed by a non-empty sequence of decimal digits. In particular, you may not put commas, periods, or spaces in the middle of large integers, as we sometimes do in English and other natural languages (e.g., ‘2,147,483,647’).

Real numbers in MOO are represented as they are in almost all other programming languages, using so-called floating-point numbers. These have certain (large) limits on size and precision that make them useful for a wide range of applications. Floating-point numbers are written with an optional minus sign followed by a non-empty sequence of digits punctuated at some point with a decimal point (‘.’) and/or followed by a scientific-notation marker (the letter ‘E’ or ‘e’ followed by an optional sign and one or more digits). Here are some examples of floating-point numbers:

     325.0   325.   3.25e2   0.325E3   325.E1   .0325e+4   32500e-2

All of these examples mean the same number. The third of these, as an example of scientific notation, should be read “3.25 times 10 to the power of 2”.

Fine points: The MOO represents floating-point numbers using the local meaning of the C-language double type, which is almost always equivalent to IEEE 754 double precision floating point. If so, then the smallest positive floating-point number is no larger than 2.2250738585072014e-308 and the largest floating-point number is

     1.7976931348623157e+308.

IEEE infinities and NaN values are not allowed in MOO. The error E_FLOAT is raised whenever an infinity would otherwise be computed; E_INVARG is raised whenever a NaN would otherwise arise. The value 0.0 is always returned on underflow.

Character strings are arbitrarily-long sequences of normal, ASCII printing characters. When written as values in a program, strings are enclosed in double-quotes, like this:

"This is a character string."

To include a double-quote in the string, precede it with a backslash (‘\’), like this:

"His name was \"Leroy\", but nobody ever called him that."

Finally, to include a backslash in a string, double it:

"Some people use backslash ('\\') to mean set difference."

MOO strings may not include special ASCII characters like carriage-return, line-feed, bell, etc. The only non-printing characters allowed are spaces and tabs.

Fine point: There is a special kind of string used for representing the arbitrary bytes used in general, binary input and output. In a binary string, any byte that isn’t an ASCII printing character or the space character is represented as the three-character substring "~XX", where XX is the hexadecimal representation of the byte; the input character ‘~’ is represented by the three-character substring "~7E". This special representation is used by the functions encode_binary() and decode_binary() and by the functions notify() and read() with network connections that are in binary

mode. See the descriptions of the set_connection_option(), encode_binary(), and

decode_binary() functions for more details. Objects are the backbone of the MOO database and, as such, deserve a great deal of discussion;

the entire next section is devoted to them. For now, let it suffice to say that every object has a number, unique to that object. In programs, we write a reference to a particular object by putting a hash mark (‘#’) followed by the number, like this:

#495

Object numbers are always integers.

There are three special object numbers used for a variety of purposes: #-1, #-2, and #-3, usually referred to in the LambdaCore database as $nothing, $ambiguous_match, and $failed_match, respectively.

Errors are, by far, the least frequently used values in MOO. In the normal case, when a program attempts an operation that is erroneous for some reason (for example, trying to add a number to a character string), the server stops running the program and prints out an error message. However, it is possible for a program to stipulate that such errors should not stop execution; instead, the server should just let the value of the operation be an error value. The program can then test for such a result and take some appropriate kind of recovery action. In programs, error values are written as words beginning with ‘E_’. The complete list of error values, along with their associated messages, is as follows:

E_NONE No error E_TYPE Type mismatch E_DIV Division by zero E_PERM Permission denied E_PROPNF Property not found E_VERBNF Verb not found E_VARNF Variable not found E_INVIND Invalid indirection E_RECMOVE Recursive move E_MAXREC Too many verb calls E_RANGE Range error E_ARGS Incorrect number of arguments E_NACC Move refused by destination E_INVARG Invalid argument E_QUOTA Resource limit exceeded E_FLOAT Floating-point arithmetic error

The final kind of value in MOO programs is lists. A list is a sequence of arbitrary MOO values, possibly including other lists. In programs, lists are written in mathematical set notation with each of the elements written out in order, separated by commas, the whole enclosed in curly braces (‘{’ and ‘}’). For example, a list of the names of the days of the week is written like this:

     {"Sunday", "Monday", "Tuesday", "Wednesday",
      "Thursday", "Friday", "Saturday"}

Note that it doesn’t matter that we put a line-break in the middle of the list. This is true in general in MOO: anywhere that a space can go, a line-break can go, with the same meaning. The only exception is inside character strings, where line-breaks are not allowed.

2.2 Objects in the MOO Database

Objects are, in a sense, the whole point of the MOO programming language. They are used to represent objects in the virtual reality, like people, rooms, exits, and other concrete things. Because of this, MOO makes a bigger deal out of creating objects than it does for other kinds of value, like integers.

Numbers always exist, in a sense; you have only to write them down in order to operate on them. With objects, it is different. The object with number ‘#958’ does not exist just because you write down its number. An explicit operation, the ‘create()’ function described later, is required to bring an object into existence. Symmetrically, once created, objects continue to exist until they are explicitly destroyed by the ‘recycle()’ function (also described later).

The identifying number associated with an object is unique to that object. It was assigned when the object was created and will never be reused, even if the object is destroyed. Thus, if we create an object and it is assigned the number ‘#1076’, the next object to be created will be assigned ‘#1077’, even if ‘#1076’ is destroyed in the meantime.

Every object is made up of three kinds of pieces that together define its behavior: attributes, properties, and verbs.

2.2.1 Fundamental Object Attributes

There are three fundamental attributes to every object:

1. A flag (either true or false) specifying whether or not the object represents a player,

2. The object that is its parent, and

3. A list of the objects that are its children; that is, those objects for which this object is their parent.

The act of creating a character sets the player attribute of an object and only a wizard (using the function set_player_flag()) can change that setting. Only characters have the player bit set to 1.

The parent/child hierarchy is used for classifying objects into general classes and then sharing behavior among all members of that class. For example, the LambdaCore database contains an object representing a sort of “generic” room. All other rooms are descendants (i.e., children or children’s children, or …) of that one. The generic room defines those pieces of behavior that are common to all rooms; other rooms specialize that behavior for their own purposes. The notion of classes and specialization is the very essence of what is meant by object-oriented programming. Only the functions create(), recycle(), chparent(), and renumber() can change the parent and children attributes.

2.2.2 Properties on Objects

A property is a named “slot” in an object that can hold an arbitrary MOO value. Every object has eight built-in properties whose values are constrained to be of particular types. In addition, an object can have any number of other properties, none of which have type constraints. The built-in properties are as follows:

name a string, the usual name for this object owner an object, the player who controls access to it location an object, where the object is in virtual reality contents a list of objects, the inverse of ‘location’ programmer a bit, does the object have programmer rights? wizard a bit, does the object have wizard rights? r a bit, is the object publicly readable? w a bit, is the object publicly writable? f a bit, is the object fertile?

The ‘name’ property is used to identify the object in various printed messages. It can only be set by a wizard or by the owner of the object. For player objects, the ‘name’ property can only be set by a wizard; this allows the wizards, for example, to check that no two players have the same name.

The ‘owner’ identifies the object that has owner rights to this object, allowing them, for example, to change the ‘name’ property. Only a wizard can change the value of this property.

The ‘location’ and ‘contents’ properties describe a hierarchy of object containment in the virtual reality. Most objects are located “inside” some other object and that other object is the value of the ‘location’ property. The ‘contents’ property is a list of those objects for which this object is their location. In order to maintain the consistency of these properties, only the move() function is able to change them.

The ‘wizard’ and ‘programmer’ bits are only applicable to characters, objects representing players. They control permission to use certain facilities in the server. They may only be set by a wizard.

The ‘r’ bit controls whether or not players other than the owner of this object can obtain a list of the properties or verbs in the object. Symmetrically, the ‘w’ bit controls whether or not non-owners can add or delete properties and/or verbs on this object. The ‘r’ and ‘w’ bits can only be set by a wizard or by the owner of the object.

The ‘f’ bit specifies whether or not this object is fertile, whether or not players other than the owner of this object can create new objects with this one as the parent. It also controls whether or not non-owners can use the chparent() built-in function to make this object the parent of an existing object. The ‘f’ bit can only be set by a wizard or by the owner of the object.

All of the built-in properties on any object can, by default, be read by any player. It is possible, however, to override this behavior from within the database, making any of these properties readable only by wizards. See the chapter on server assumptions about the database for details.

As mentioned above, it is possible, and very useful, for objects to have other properties aside from the built-in ones. These can come from two sources.

First, an object has a property corresponding to every property in its parent object. To use the jargon of object-oriented programming, this is a kind of inheritance. If some object has a property named ‘foo’, then so will all of its children and thus its children’s children, and so on.

Second, an object may have a new property defined only on itself and its descendants. For example, an object representing a rock might have properties indicating its weight, chemical composition, and/or pointiness, depending upon the uses to which the rock was to be put in the virtual reality.

Every defined property (as opposed to those that are built-in) has an owner and a set of permissions for non-owners. The owner of the property can get and set the property’s value and can change the non-owner permissions. Only a wizard can change the owner of a property.

The initial owner of a property is the player who added it; this is usually, but not always, the player who owns the object to which the property was added. This is because properties can only be added by the object owner or a wizard, unless the object is publicly writable (i.e., its ‘w’ property is 1), which is rare. Thus, the owner of an object may not necessarily be the owner of every (or even any) property on that object.

The permissions on properties are drawn from this set: ‘r’ (read), ‘w’ (write), and ‘c’ (change ownership in descendants). Read permission lets non-owners get the value of the property and, of course, write permission lets them set that value. The ‘c’ permission bit is a little more complicated.

Recall that every object has all of the properties that its parent does and perhaps some more. Ordinarily, when a child object inherits a property from its parent, the owner of the child becomes the owner of that property. This is because the ‘c’ permission bit is “on” by default. If the ‘c’ bit is not on, then the inherited property has the same owner in the child as it does in the parent.

As an example of where this can be useful, the LambdaCore database ensures that every player has a ‘password’ property containing the encrypted version of the player’s connection password. For security reasons, we don’t want other players to be able to see even the encrypted version of the password, so we turn off the ‘r’ permission bit. To ensure that the password is only set in a consistent way (i.e., to the encrypted version of a player’s password), we don’t want to let anyone but a wizard change the property. Thus, in the parent object for all players, we made a wizard the owner of the password property and set the permissions to the empty string, "". That is, non-owners cannot read or write the property and, because the ‘c’ bit is not set, the wizard who owns the property on the parent class also owns it on all of the descendants of that class.

Another, perhaps more down-to-earth example arose when a character named Ford started building objects he called “radios” and another character, yduJ, wanted to own one. Ford kindly made the generic radio object fertile, allowing yduJ to create a child object of it, her own radio. Radios had a property called ‘channel’ that identified something corresponding to the frequency to which the radio was tuned. Ford had written nice programs on radios (verbs, discussed below) for turning the channel selector on the front of the radio, which would make a corresponding change in the value of the ‘channel’ property. However, whenever anyone tried to turn the channel selector on yduJ’s radio, they got a permissions error. The problem concerned the ownership of the ‘channel’ property.

As I explain later, programs run with the permissions of their author. So, in this case, Ford’s nice verb for setting the channel ran with his permissions. But, since the ‘channel’ property in the generic radio had the ‘c’ permission bit set, the ‘channel’ property on yduJ’s radio was owned by her. Ford didn’t have permission to change it! The fix was simple. Ford changed the permissions on the ‘channel’ property of the generic radio to be just ‘r’, without the ‘c’ bit, and yduJ made a new radio. This time, when yduJ’s radio inherited the ‘channel’ property, yduJ did not inherit ownership of it; Ford remained the owner. Now the radio worked properly, because Ford’s verb had permission to change the channel.

2.2.3 Verbs on Objects

The final kind of piece making up an object is verbs. A verb is a named MOO program that is associated with a particular object. Most verbs implement commands that a player might type; for example, in the LambdaCore database, there is a verb on all objects representing containers that implements commands of the form ‘put object in container’. It is also possible for MOO programs to invoke the verbs defined on objects. Some verbs, in fact, are designed to be used only from inside MOO code; they do not correspond to any particular player command at all. Thus, verbs in MOO are like the ‘procedures’ or ‘methods’ found in some other programming languages.

As with properties, every verb has an owner and a set of permission bits. The owner of a verb can change its program, its permission bits, and its argument specifiers (discussed below). Only a wizard can change the owner of a verb. The owner of a verb also determines the permissions with which that verb runs; that is, the program in a verb can do whatever operations the owner of that verb is allowed to do and no others. Thus, for example, a verb owned by a wizard must be written very carefully, since wizards are allowed to do just about anything.

The permission bits on verbs are drawn from this set: ‘r’ (read), ‘w’ (write), ‘x’ (execute), and ‘d’ (debug). Read permission lets non-owners see the program for a verb and, symmetrically, write permission lets them change that program. The other two bits are not, properly speaking, permission bits at all; they have a universal effect, covering both the owner and non-owners.

The execute bit determines whether or not the verb can be invoked from within a MOO program (as opposed to from the command line, like the ‘put’ verb on containers). If the ‘x’ bit is not set, the verb cannot be called from inside a program. The ‘x’ bit is usually set.

The setting of the debug bit determines what happens when the verb’s program does something erroneous, like subtracting a number from a character string. If the ‘d’ bit is set, then the server raises an error value; such raised errors can be caught by certain other pieces of MOO code. If the error is not caught, however, the server aborts execution of the command and, by default, prints an error message on the terminal of the player whose command is being executed. (See the chapter on server assumptions about the database for details on how uncaught errors are handled.) If the ‘d’ bit is not set, then no error is raised, no message is printed, and the command is not aborted; instead the error value is returned as the result of the erroneous operation.

Note: the ‘d’ bit exists only for historical reasons; it used to be the only way for MOO code to catch and handle errors. With the introduction of the try-except statement and the error-catching expression, the ‘d’ bit is no longer useful. All new verbs should have the ‘d’ bit set, using the newer facilities for error handling if desired. Over time, old verbs written assuming the ‘d’ bit would not be set should be changed to use the new facilities instead.

In addition to an owner and some permission bits, every verb has three ‘argument specifiers’, one each for the direct object, the preposition, and the indirect object. The direct and indirect specifiers are each drawn from this set: ‘this’, ‘any’, or ‘none’. The preposition specifier is ‘none’, ‘any’, or one of the items in this list:

with/using
at/to
in front of
in/inside/into
on top of/on/onto/upon
out of/from inside/from
over
through
under/underneath/beneath
behind
beside
for/about
is
as
off/off of

The argument specifiers are used in the process of parsing commands, described in the next chapter.

3 The Built-in Command Parser

The MOO server is able to do a small amount of parsing on the commands that a player enters. In particular, it can break apart commands that follow one of the following forms:

verb

verb direct-object

verb direct-object preposition indirect-object Real examples of these forms, meaningful in the LambdaCore database, are as follows:

     look
     take yellow bird
     put yellow bird in cuckoo clock

Note that English articles (i.e., ‘the’, ‘a’, and ‘an’) are not generally used in MOO commands; the parser does not know that they are not important parts of objects’ names.

To have any of this make real sense, it is important to understand precisely how the server decides what to do when a player types a command.

First, the server checks whether or not the first non-blank character in the command is one of the following:

     "        :        ;

If so, that character is replaced by the corresponding command below, followed by a space:

     say      emote    eval

For example, the command

     "Hi, there.

is treated exactly as if it were as follows:

     say Hi, there.

The server next breaks up the command into words. In the simplest case, the command is broken into words at every run of space characters; for example, the command ‘foo bar baz’ would be broken into the words ‘foo’, ‘bar’, and ‘baz’. To force the server to include spaces in a “word”, all or part of a word can be enclosed in double-quotes. For example, the command

     foo "bar mumble" baz" "fr"otz" bl"o"rt

is broken into the words ‘foo’, ‘bar mumble’, ‘baz frotz’, and ‘blort’. Finally, to include a doublequote or a backslash in a word, they can be preceded by a backslash, just like in MOO strings.

Having thus broken the string into words, the server next checks to see if the first word names any

of the six “built-in” commands: ‘.program’, ‘PREFIX’, ‘OUTPUTPREFIX’, ‘SUFFIX’, ‘OUTPUTSUFFIX’,

or the connection’s defined flush command, if any (‘.flush’ by default). The first one of these is only available to programmers, the next four are intended for use by client programs, and the last can vary from database to database or even connection to connection; all six are described in the final chapter of this document, “Server Commands and Database Assumptions”. If the first word isn’t one of the above, then we get to the usual case: a normal MOO command.

The server next gives code in the database a chance to handle the command. If the verb $do_ command() exists, it is called with the words of the command passed as its arguments and argstr set to the raw command typed by the user. If $do_command() does not exist, or if that verb-call completes normally (i.e., without suspending or aborting) and returns a false value, then the builtin command parser is invoked to handle the command as described below. Otherwise, it is assumed that the database code handled the command completely and no further action is taken by the server for that command.

If the built-in command parser is invoked, the server tries to parse the command into a verb, direct object, preposition and indirect object. The first word is taken to be the verb. The server then tries to find one of the prepositional phrases listed at the end of the previous section, using the match that occurs earliest in the command. For example, in the very odd command ‘foo as bar to baz’, the server would take ‘as’ as the preposition, not ‘to’.

If the server succeeds in finding a preposition, it considers the words between the verb and the preposition to be the direct object and those after the preposition to be the indirect object. In both cases, the sequence of words is turned into a string by putting one space between each pair of words. Thus, in the odd command from the previous paragraph, there are no words in the direct object (i.e., it is considered to be the empty string, "") and the indirect object is "bar to baz".

If there was no preposition, then the direct object is taken to be all of the words after the verb and the indirect object is the empty string.

The next step is to try to find MOO objects that are named by the direct and indirect object strings.

First, if an object string is empty, then the corresponding object is the special object #-1 (aka $nothing in LambdaCore). If an object string has the form of an object number (i.e., a hash mark (‘#’) followed by digits), and the object with that number exists, then that is the named object. If the object string is either "me" or "here", then the player object itself or its location is used, respectively.

Otherwise, the server considers all of the objects whose location is either the player (i.e., the objects the player is “holding”, so to speak) or the room the player is in (i.e., the objects in the same room as the player); it will try to match the object string against the various names for these objects.

The matching done by the server uses the ‘aliases’ property of each of the objects it considers. The value of this property should be a list of strings, the various alternatives for naming the object. If it is not a list, or the object does not have an ‘aliases’ property, then the empty list is used. In any case, the value of the ‘name’ property is added to the list for the purposes of matching.

The server checks to see if the object string in the command is either exactly equal to or a prefix of any alias; if there are any exact matches, the prefix matches are ignored. If exactly one of the objects being considered has a matching alias, that object is used. If more than one has a match, then the special object #-2 (aka $ambiguous_match in LambdaCore) is used. If there are no matches, then the special object #-3 (aka $failed_match in LambdaCore) is used.

So, now the server has identified a verb string, a preposition string, and direct- and indirectobject strings and objects. It then looks at each of the verbs defined on each of the following four objects, in order:

1. the player who typed the command,

2. the room the player is in,

3. the direct object, if any, and

4. the indirect object, if any.

For each of these verbs in turn, it tests if all of the the following are true:

• the verb string in the command matches one of the names for the verb, • the direct- and indirect-object values found by matching are allowed by the corresponding

argument specifiers for the verb, and

• the preposition string in the command is matched by the preposition specifier for the verb.

I’ll explain each of these criteria in turn.

Every verb has one or more names; all of the names are kept in a single string, separated by spaces. In the simplest case, a verb-name is just a word made up of any characters other than spaces and stars (i.e., ‘ ’ and ‘*’). In this case, the verb-name matches only itself; that is, the name must be matched exactly.

If the name contains a single star, however, then the name matches any prefix of itself that is at least as long as the part before the star. For example, the verb-name ‘foo*bar’ matches any of the strings ‘foo’, ‘foob’, ‘fooba’, or ‘foobar’; note that the star itself is not considered part of the name.

If the verb name ends in a star, then it matches any string that begins with the part before the star. For example, the verb-name ‘foo*’ matches any of the strings ‘foo’, ‘foobar’, ‘food’, or ‘foogleman’, among many others. As a special case, if the verb-name is ‘*’ (i.e., a single star all by itself), then it matches anything at all.

Recall that the argument specifiers for the direct and indirect objects are drawn from the set ‘none’, ‘any’, and ‘this’. If the specifier is ‘none’, then the corresponding object value must be #-1 (aka $nothing in LambdaCore); that is, it must not have been specified. If the specifier is ‘any’, then the corresponding object value may be anything at all. Finally, if the specifier is ‘this’, then the corresponding object value must be the same as the object on which we found this verb; for example, if we are considering verbs on the player, then the object value must be the player object.

Finally, recall that the argument specifier for the preposition is either ‘none’, ‘any’, or one of several sets of prepositional phrases, given above. A specifier of ‘none’ matches only if there was no preposition found in the command. A specifier of ‘any’ always matches, regardless of what preposition was found, if any. If the specifier is a set of prepositional phrases, then the one found must be in that set for the specifier to match.

So, the server considers several objects in turn, checking each of their verbs in turn, looking for the first one that meets all of the criteria just explained. If it finds one, then that is the verb whose program will be executed for this command. If not, then it looks for a verb named ‘huh’ on the room that the player is in; if one is found, then that verb will be called. This feature is useful for implementing room-specific command parsing or error recovery. If the server can’t even find a ‘huh’ verb to run, it prints an error message like ‘I couldn't understand that.’ and the command is considered complete.

At long last, we have a program to run in response to the command typed by the player. When the code for the program begins execution, the following built-in variables will have the indicated values:

player an object, the player who typed the command this an object, the object on which this verb was found caller an object, the same as ‘player’ verb a string, the first word of the command argstr a string, everything after the first word of the command args a list of strings, the words in ‘argstr’ dobjstr a string, the direct object string found during parsing dobj an object, the direct object value found during matching prepstr a string, the prepositional phrase found during parsing iobjstr a string, the indirect object string iobj an object, the indirect object value

The value returned by the program, if any, is ignored by the server.

4 The MOO Programming Language

MOO stands for “MUD, Object Oriented.” MUD, in turn, has been said to stand for many different things, but I tend to think of it as “Multi-User Dungeon” in the spirit of those ancient precursors to MUDs, Adventure and Zork.

MOO, the programming language, is a relatively small and simple object-oriented language designed to be easy to learn for most non-programmers; most complex systems still require some significant programming ability to accomplish, however.

Having given you enough context to allow you to understand exactly what MOO code is doing, I now explain what MOO code looks like and what it means. I begin with the syntax and semantics of expressions, those pieces of code that have values. After that, I cover statements, the next level of structure up from expressions. Next, I discuss the concept of a task, the kind of running process initiated by players entering commands, among other causes. Finally, I list all of the built-in functions available to MOO code and describe what they do.

First, though, let me mention comments. You can include bits of text in your MOO program that are ignored by the server. The idea is to allow you to put in notes to yourself and others about what the code is doing. To do this, begin the text of the comment with the two characters ‘/*’ and end it with the two characters ‘*/’; this is just like comments in the C programming language. Note that the server will completely ignore that text; it will not be saved in the database. Thus, such comments are only useful in files of code that you maintain outside the database.

To include a more persistent comment in your code, try using a character string literal as a statement. For example, the sentence about peanut butter in the following code is essentially ignored during execution but will be maintained in the database:

     for x in (players())
       "Grendel eats peanut butter!";
       player:tell(x.name, " (", x, ")");
     endfor

4.1 MOO Language Expressions

Expressions are those pieces of MOO code that generate values; for example, the MOO code

     3 + 4

is an expression that generates (or “has” or “returns”) the value 7. There are many kinds of expressions in MOO, all of them discussed below.

4.1.1 Errors While Evaluating Expressions

Most kinds of expressions can, under some circumstances, cause an error to be generated. For example, the expression x / y will generate the error E_DIV if y is equal to zero. When an expression generates an error, the behavior of the server is controlled by setting of the ‘d’ (debug) bit on the verb containing that expression. If the ‘d’ bit is not set, then the error is effectively squelched immediately upon generation; the error value is simply returned as the value of the expression that generated it.

Note: this error-squelching behavior is very error prone, since it affects all errors, including ones the programmer may not have anticipated. The ‘d’ bit exists only for historical reasons; it was once the only way for MOO programmers to catch and handle errors. The error-catching expression and the try-except statement, both described below, are far better ways of accomplishing the same thing.

If the ‘d’ bit is set, as it usually is, then the error is raised and can be caught and handled either by code surrounding the expression in question or by verbs higher up on the chain of calls leading to the current verb. If the error is not caught, then the server aborts the entire task and, by default, prints a message to the current player. See the descriptions of the error-catching expression and the try-except statement for the details of how errors can be caught, and the chapter on server assumptions about the database for details on the handling of uncaught errors.

4.1.2 Writing Values Directly in Verbs

The simplest kind of expression is a literal MOO value, just as described in the section on values at the beginning of this document. For example, the following are all expressions:

     17
     #893
     "This is a character string."
     E_TYPE
     {"This", "is", "a", "list", "of", "words"}

In the case of lists, like the last example above, note that the list expression contains other expressions, several character strings in this case. In general, those expressions can be of any kind at all, not necessarily literal values. For example,

     {3 + 4, 3 - 4, 3 * 4}

is an expression whose value is the list {7, -1, 12}.

4.1.3 Naming Values Within a Verb

As discussed earlier, it is possible to store values in properties on objects; the properties will keep those values forever, or until another value is explicitly put there. Quite often, though, it is useful to have a place to put a value for just a little while. MOO provides local variables for this purpose.

Variables are named places to hold values; you can get and set the value in a given variable as many times as you like. Variables are temporary, though; they only last while a particular verb is running; after it finishes, all of the variables given values there cease to exist and the values are forgotten.

Variables are also “local” to a particular verb; every verb has its own set of them. Thus, the variables set in one verb are not visible to the code of other verbs.

The name for a variable is made up entirely of letters, digits, and the underscore character (‘_’) and does not begin with a digit. The following are all valid variable names:

     foo
     _foo
     this2that
     M68000
     two_words
     This_is_a_very_long_multiword_variable_name

Note that, along with almost everything else in MOO, the case of the letters in variable names is insignificant. For example, these are all names for the same variable:

     fubar
     Fubar
     FUBAR
     fUbAr

A variable name is itself an expression; its value is the value of the named variable. When a verb begins, almost no variables have values yet; if you try to use the value of a variable that doesn’t have one, the error value E_VARNF is raised. (MOO is unlike many other programming languages in which one must ‘declare’ each variable before using it; MOO has no such declarations.) The following variables always have values:

     INT         FLOAT        OBJ
     STR         LIST         ERR
     player      this         caller
     verb        args         argstr
     dobj        dobjstr      prepstr
     iobj        iobjstr      NUM

The values of some of these variables always start out the same:

INT an integer, the type code for integers (see the description of the function typeof(), below)

NUM the same as INT (for historical reasons)

FLOAT an integer, the type code for floating-point numbers

LIST an integer, the type code for lists

STR an integer, the type code for strings

OBJ an integer, the type code for objects

ERR an integer, the type code for error values

For others, the general meaning of the value is consistent, though the value itself is different for different situations:

player an object, the player who typed the command that started the task that involved running this piece of code.

this an object, the object on which the currently-running verb was found.

caller an object, the object on which the verb that called the currently-running verb was found. For the first verb called for a given command, ‘caller’ has the same value as

          ‘player’.

verb a string, the name by which the currently-running verb was identified.

args a list, the arguments given to this verb. For the first verb called for a given command, this is a list of strings, the words on the command line.

The rest of the so-called “built-in” variables are only really meaningful for the first verb called for a given command. Their semantics is given in the discussion of command parsing, above.

To change what value is stored in a variable, use an assignment expression:

variable = expression For example, to change the variable named ‘x’ to have the value 17, you would write ‘x = 17’ as

an expression. An assignment expression does two things:

• it changes the value of of the named variable, and

• it returns the new value of that variable.

Thus, the expression

13 + (x = 17)

changes the value of ‘x’ to be 17 and returns 30.

4.1.4 Arithmetic Operators

All of the usual simple operations on numbers are available to MOO programs:

+    -    *    /    %

These are, in order, addition, subtraction, multiplication, division, and remainder. In the following table, the expressions on the left have the corresponding values on the right:

5 + 2  ⇒ 7
5 - 2  ⇒ 3
5 * 2  ⇒ 10
5 / 2  ⇒ 2
5.0 / 2.0 ⇒ 2.5
5 % 2  ⇒ 1
5.0 % 2.0 ⇒ 1.0
5 % -2  ⇒ 1
-5 % 2  ⇒ -1
-5 % -2  ⇒ -1
-(5 + 2)  ⇒ -7

Note that integer division in MOO throws away the remainder and that the result of the remainder operator (‘%’) has the same sign as the left-hand operand. Also, note that ‘-’ can be used without a left-hand operand to negate a numeric expression.

Fine point: Integers and floating-point numbers cannot be mixed in any particular use of these arithmetic operators; unlike some other programming languages, MOO does not automatically coerce integers into floating-point numbers. You can use the tofloat() function to perform an explicit conversion.

The ‘+’ operator can also be used to append two strings. The expression

"foo" + "bar"

has the value

"foobar"

Unless both operands to an arithmetic operator are numbers of the same kind (or, for ‘+’, both strings), the error value E_TYPE is raised. If the right-hand operand for the division or remainder operators (‘/’ or ‘%’) is zero, the error value E_DIV is raised.

MOO also supports the exponentiation operation, also known as “raising to a power,” using the ‘^’ operator:

3 ^4  ⇒ 81
3 ^4.5      error  E_TYPE
3.5 ^4  ⇒ 150.0625
3.5 ^4.5 ⇒ 280.741230801382

Note that if the first operand is an integer, then the second operand must also be an integer. If the first operand is a floating-point number, then the second operand can be either kind of number. Although it is legal to raise an integer to a negative power, it is unlikely to be terribly useful.

4.1.5 Comparing Values

Any two values can be compared for equality using ‘==’ and ‘!=’. The first of these returns 1 if the two values are equal and 0 otherwise; the second does the reverse:

3 == 4  ⇒ 0
3 != 4  ⇒ 1
3 == 3.0  ⇒ 0
"foo" == "Foo"  ⇒ 1
#34 != #34  ⇒ 0
{1, #34, "foo"} == {1, #34, "FoO"} ⇒ 1
E_DIV == E_TYPE  ⇒ 0
3 != "foo"  ⇒ 1

Note that integers and floating-point numbers are never equal to one another, even in the ‘obvious’ cases. Also note that comparison of strings (and list values containing strings) is case-insensitive; that is, it does not distinguish between the upper- and lower-case version of letters. To test two values for case-sensitive equality, use the ‘equal’ function described later.

Warning: It is easy (and very annoying) to confuse the equality-testing operator (‘==’) with the assignment operator (‘=’), leading to nasty, hard-to-find bugs. Don’t do this.

Numbers, object numbers, strings, and error values can also be compared for ordering purposes using the following operators:

<       <=      >=      >

meaning “less than,” “less than or equal,” “greater than or equal,” and “greater than,” respectively. As with the equality operators, these return 1 when their operands are in the appropriate relation and 0 otherwise:

3 < 4  ⇒ 1
3 < 4.0          error E_TYPE
#34 >= #32  ⇒ 1
"foo" <= "Boo" ⇒ 0
E_DIV > E_TYPE ⇒ 1

Note that, as with the equality operators, strings are compared case-insensitively. To perform a case-sensitive string comparison, use the ‘strcmp’ function described later. Also note that the error values are ordered as given in the table in the section on values. If the operands to these four comparison operators are of different types (even integers and floating-point numbers are considered different types), or if they are lists, then E_TYPE is raised.

4.1.6 Values as True and False

There is a notion in MOO of true and false values; every value is one or the other. The true values are as follows:

• all integers other than zero,

• all floating-point numbers not equal to 0.0,

• all non-empty strings (i.e., other than ‘""’), and

• all non-empty lists (i.e., other than ‘{}’).

All other values are false:

• the integer zero,

• the floating-point numbers 0.0 and -0.0,

• the empty string (‘""’),

• the empty list (‘{}’),

• all object numbers, and

• all error values.

There are four kinds of expressions and two kinds of statements that depend upon this classification of MOO values. In describing them, I sometimes refer to the truth value of a MOO value; this is just true or false, the category into which that MOO value is classified.

The conditional expression in MOO has the following form:

expression-1 ? expression-2 | expression-3 First, expression-1 is evaluated. If it returns a true value, then expression-2 is evaluated and

whatever it returns is returned as the value of the conditional expression as a whole. If expression-1 returns a false value, then expression-3 is evaluated instead and its value is used as that of the conditional expression.

     1 ? 2 | 3  ⇒ 2

     0 ? 2 | 3  ⇒ 3
     "foo" ? 17 | {#34} ⇒ 17

Note that only one of expression-2 and expression-3 is evaluated, never both.

To negate the truth value of a MOO value, use the ‘!’ operator:

! expression If the value of expression is true, ‘!’ returns 0; otherwise, it returns 1:

     ! "foo"  ⇒ 0
     ! (3 >= 4) ⇒ 1

The negation operator is usually read as “not.”

It is frequently useful to test more than one condition to see if some or all of them are true. MOO provides two operators for this:

expression-1 && expression-2 expression-1 || expression-2

These operators are usually read as “and” and “or,” respectively.

The ‘&&’ operator first evaluates expression-1. If it returns a true value, then expression-2 is evaluated and its value becomes the value of the ‘&&’ expression as a whole; otherwise, the value of expression-1 is used as the value of the ‘&&’ expression. Note that expression-2 is only evaluated if expression-1 returns a true value. The ‘&&’ expression is equivalent to the conditional expression

expression-1 ? expression-2 | expression-1

except that expression-1 is only evaluated once.

The ‘||’ operator works similarly, except that expression-2 is evaluated only if expression-1 returns a false value. It is equivalent to the conditional expression

expression-1 ? expression-1 | expression-2

except that, as with ‘&&’, expression-1 is only evaluated once.

These two operators behave very much like “and” and “or” in English:

     1 && 1  ⇒ 1
     0 && 1  ⇒ 0
     0 && 0  ⇒ 0
     1 || 1  ⇒ 1
     0 || 1  ⇒ 1
     0 || 0  ⇒ 0
     17 <= 23  &&  23 <= 27 ⇒ 1

4.1.7 Indexing into Lists and Strings

Both strings and lists can be seen as ordered sequences of MOO values. In the case of strings, each is a sequence of single-character strings; that is, one can view the string "bar" as a sequence of the strings "b", "a", and "r". MOO allows you to refer to the elements of lists and strings by number, by the index of that element in the list or string. The first element in a list or string has index 1, the second has index 2, and so on.

4.1.7.1 Extracting an Element from a List or String

The indexing expression in MOO extracts a specified element from a list or string:

expression-1[expression-2]

First, expression-1 is evaluated; it must return a list or a string (the sequence). Then, expression-2 is evaluated and must return an integer (the index). If either of the expressions returns some other type of value, E_TYPE is returned. The index must be between 1 and the length of the sequence, inclusive; if it is not, then E_RANGE is raised. The value of the indexing expression is the index’th element in the sequence. Anywhere within expression-2, you can use the symbol $ as an expression returning the length of the value of expression-1.

     "fob"[2]  ⇒ "o"
     "fob"[1]  ⇒ "f"
     {#12, #23, #34}[$ - 1] ⇒ #23

Note that there are no legal indices for the empty string or list, since there are no integers between 1 and 0 (the length of the empty string or list).

Fine point: The $ expression actually returns the length of the value of the expression just before the nearest enclosing […] indexing or subranging brackets. For example:

          "frob"[{3, 2, 4}[$]]  ⇒ "b"

4.1.7.2 Replacing an Element of a List or String

It often happens that one wants to change just one particular slot of a list or string, which is stored in a variable or a property. This can be done conveniently using an indexed assignment having one of the following forms:

variable[index-expr] = result-expr

object-expr.name[index-expr] = result-expr

object-expr.(name-expr)[index-expr] = result-expr

$name[index-expr] = result-expr The first form writes into a variable, and the last three forms write into a property. The usual errors (E_TYPE, E_INVIND, E_PROPNF and E_PERM for lack of read/write permission on the property) may be raised, just as in reading and writing any object property; see the discussion of object property expressions below for details. Correspondingly, if variable does not yet have a value (i.e., it has never been assigned to), E_VARNF will be raised.

If index-expr is not an integer, or if the value of variable or the property is not a list or string, E_TYPE is raised. If result-expr is a string, but not of length 1, E_INVARG is raised. Now suppose index-expr evaluates to an integer k. If k is outside the range of the list or string (i.e. smaller than 1 or greater than the length of the list or string), E_RANGE is raised. Otherwise, the actual assignment takes place. For lists, the variable or the property is assigned a new list that is identical to the original one except at the k-th position, where the new list contains the result of result-expr instead. For strings, the variable or the property is assigned a new string that is identical to the original one, except the k-th character is changed to be result-expr.

The assignment expression itself returns the value of result-expr. For the following examples, assume that l initially contains the list {1, 2, 3} and that s initially contains the string "foobar":

     l[5] = 3           error  E_RANGE
     l["first"] = 4     error  E_TYPE
     s[3] = "baz"       error  E_INVARG
     l[2] = l[2] + 3 ⇒ 5
     l  ⇒ {1, 5, 3}
     l[2] = "foo"  ⇒ "foo"
     l  ⇒ {1, "foo", 3}
     s[2] = "u"  ⇒ "u"
     s  ⇒ "fuobar"
     s[$] = "z"  ⇒ "z"
     s  ⇒ "fuobaz"

Note that the $ expression may also be used in indexed assignments with the same meaning as before.

Fine point: After an indexed assignment, the variable or property contains a new list or string, a copy of the original list in all but the k-th place, where it contains a new value. In programming-language jargon, the original list is not mutated, and there is no aliasing. (Indeed, no MOO value is mutable and no aliasing ever occurs.)

In the list case, indexed assignment can be nested to many levels, to work on nested lists. Assume that l initially contains the list

     {{1, 2, 3}, {4, 5, 6}, "foo"}

in the following examples:

l[7] = 4              error  E_RANGE
l[1][8] = 35          error  E_RANGE
l[3][2] = 7           error  E_TYPE
l[1][1][1] = 3        error  E_TYPE
l[2][2] = -l[2][2] ⇒ -5
l  ⇒ {{1, 2, 3}, {4, -5, 6}, "foo"}
l[2] = "bar"  ⇒ "bar"
l  ⇒ {{1, 2, 3}, "bar", "foo"}
l[2][$] = "z"  ⇒ "z"
l  ⇒ {{1, 2, 3}, "baz", "foo"}

The first two examples raise E_RANGE because 7 is out of the range of l and 8 is out of the range of l[1]. The next two examples raise E_TYPE because l[3] and l[1][1] are not lists.

4.1.7.3 Extracting a Subsequence of a List or String

The range expression extracts a specified subsequence from a list or string:

expression-1[expression-2..expression-3] The three expressions are evaluated in order. Expression-1 must return a list or string (the

sequence) and the other two expressions must return integers (the low and high indices, respectively); otherwise, E_TYPE is raised. The $ expression can be used in either or both of expression-2 and expression-3 just as before, meaning the length of the value of expression-1.

If the low index is greater than the high index, then the empty string or list is returned, depending on whether the sequence is a string or a list. Otherwise, both indices must be between 1 and the length of the sequence; E_RANGE is raised if they are not. A new list or string is returned that contains just the elements of the sequence with indices between the low and high bounds.

"foobar"[2..$]  ⇒ "oobar"
"foobar"[3..3]  ⇒ "o"
"foobar"[17..12]  ⇒ ""
{"one", "two", "three"}[$ - 1..$]  ⇒ {"two", "three"}
{"one", "two", "three"}[3..3]  ⇒ {"three"}
{"one", "two", "three"}[17..12]  ⇒ {}

4.1.7.4 Replacing a Subsequence of a List or String

The subrange assigment replaces a specified subsequence of a list or string with a supplied subsequence. The allowed forms are:

variable[start-index-expr..end-index-expr] = result-expr object-expr.name[start-index-expr..end-index-expr] = result-expr object-expr.(name-expr)[start-index-expr..end-index-expr] = result-expr $name[start-index-expr..end-index-expr] = result-expr

As with indexed assigments, the first form writes into a variable, and the last three forms write into a property. The same errors (E_TYPE, E_INVIND, E_PROPNF and E_PERM for lack of read/write permission on the property) may be raised. If variable does not yet have a value (i.e., it has never been assigned to), E_VARNF will be raised. As before, the $ expression can be used in either startindex-expr or end-index-expr, meaning the length of the original value of the expression just before the […] part.

If start-index-expr or end-index-expr is not an integer, if the value of variable or the property is not a list or string, or result-expr is not the same type as variable or the property, E_TYPE is raised. E_RANGE is raised if end-index-expr is less than zero or if start-index-expr is greater than the length of the list or string plus one. Note: the length of result-expr does not need to be the same as the length of the specified range.

In precise terms, the subrange assigment v[start..end] = value

is equivalent to

v = {@v[1..start - 1], @value, @v[end + 1..$]}

if v is a list and to

v = v[1..start - 1] + value + v[end + 1..$]

if v is a string. The assigment expression itself returns the value of result-expr. For the following examples,

assume that l initially contains the list {1, 2, 3} and that s initially contains the string "foobar":

l[5..6] = {7, 8}        error  E_RANGE
l[2..3] = 4             error  E_TYPE
l[#2..3] = {7}          error  E_TYPE
s[2..3] = {6}           error  E_TYPE
l[2..3] = {6, 7, 8, 9} ⇒ {6, 7, 8, 9}
l  ⇒ {1, 6, 7, 8, 9}
l[2..1] = {10, "foo"} ⇒ {10, "foo"}
l  ⇒ {1, 10, "foo", 6, 7, 8, 9}
l[3][2..$] = "u"  ⇒ "u"
l  ⇒ {1, 10, "fu", 6, 7, 8, 9}
s[7..12] = "baz"  ⇒ "baz"
s  ⇒ "foobarbaz"
s[1..3] = "fu"  ⇒ "fu"
s  ⇒ "fubarbaz"
s[1..0] = "test"  ⇒ "test"
s  ⇒ "testfubarbaz"

4.1.8 Other Operations on Lists

As was mentioned earlier, lists can be constructed by writing a comma-separated sequence of expressions inside curly braces:

{expression-1, expression-2, …, expression-N} The resulting list has the value of expression-1 as its first element, that of expression-2 as the second, etc.

{3 < 4, 3 <= 4, 3 >= 4, 3 > 4} ⇒ {1, 1, 0, 0}

Additionally, one may precede any of these expressions by the splicing operator, ‘@’. Such an expression must return a list; rather than the old list itself becoming an element of the new list, all of the elements of the old list are included in the new list. This concept is easy to understand, but hard to explain in words, so here are some examples. For these examples, assume that the variable a has the value {2, 3, 4} and that b has the value {"Foo", "Bar"}:

     {1, a, 5} ⇒ {1, {2, 3, 4}, 5}
     {1, @a, 5} ⇒ {1, 2, 3, 4, 5}
     {a, @a}  ⇒ {{2, 3, 4}, 2, 3, 4}
     {@a, @b}  ⇒ {2, 3, 4, "Foo", "Bar"}

If the splicing operator (‘@’) precedes an expression whose value is not a list, then E_TYPE is raised as the value of the list construction as a whole.

The list membership expression tests whether or not a given MOO value is an element of a given list and, if so, with what index:

expression-1 in expression-2 Expression-2 must return a list; otherwise, E_TYPE is raised. If the value of expression-1 is in

that list, then the index of its first occurrence in the list is returned; otherwise, the ‘in’ expression returns 0.

     2 in {5, 8, 2, 3}  ⇒ 3
     7 in {5, 8, 2, 3}  ⇒ 0
     "bar" in {"Foo", "Bar", "Baz"} ⇒ 2

Note that the list membership operator is case-insensitive in comparing strings, just like the comparison operators. To perform a case-sensitive list membership test, use the ‘is_member’ function described later. Note also that since it returns zero only if the given value is not in the given list, the ‘in’ expression can be used either as a membership test or as an element locator.

4.1.9 Spreading List Elements Among Variables

It is often the case in MOO programming that you will want to access the elements of a list individually, with each element stored in a separate variables. This desire arises, for example, at the beginning of almost every MOO verb, since the arguments to all verbs are delivered all bunched together in a single list. In such circumstances, you could write statements like these:

     first = args[1];
     second = args[2];
     if (length(args) > 2)
       third = args[3];
     else
       third = 0;
     endif

This approach gets pretty tedious, both to read and to write, and it’s prone to errors if you mistype one of the indices. Also, you often want to check whether or not any extra list elements were present, adding to the tedium.

MOO provides a special kind of assignment expression, called scattering assignment made just for cases such as these. A scattering assignment expression looks like this:

{target, …} = expr where each target describes a place to store elements of the list that results from evaluating expr. A target has one of the following forms:

variable This is the simplest target, just a simple variable; the list element in the corresponding position is assigned to the variable. This is called a required target, since the assignment is required to put one of the list elements into the variable.

?variable This is called an optional target, since it doesn’t always get assigned an element. If there are any list elements left over after all of the required targets have been accounted for (along with all of the other optionals to the left of this one), then this variable is treated like a required one and the list element in the corresponding position is assigned to the variable. If there aren’t enough elements to assign one to this target, then no assignment is made to this variable, leaving it with whatever its previous value was.

?variable = default-expr This is also an optional target, but if there aren’t enough list elements available to assign one to this target, the result of evaluating default-expr is assigned to it instead. Thus, default-expr provides a default value for the variable. The default value expressions are evaluated and assigned working from left to right after all of the other assignments have been performed.

@variable By analogy with the @ syntax in list construction, this variable is assigned a list of all of the ‘leftover’ list elements in this part of the list after all of the other targets have been filled in. It is assigned the empty list if there aren’t any elements left over. This is called a rest target, since it gets the rest of the elements. There may be at most one rest target in each scattering assignment expression.

If there aren’t enough list elements to fill all of the required targets, or if there are more than enough to fill all of the required and optional targets but there isn’t a rest target to take the leftover ones, then E_ARGS is raised.

Here are some examples of how this works. Assume first that the verb me:foo() contains the following code:

b = c = e = 17;
{a, ?b, ?c = 8, @d, ?e = 9, f} = args;
return {a, b, c, d, e, f};

Then the following calls return the given values:

me:foo(1)                         error  E_ARGS
me:foo(1, 2)  ⇒ {1, 17, 8, {}, 9, 2}
me:foo(1, 2, 3)  ⇒ {1, 2, 8, {}, 9, 3}
me:foo(1, 2, 3, 4)  ⇒ {1, 2, 3, {}, 9, 4}
me:foo(1, 2, 3, 4, 5)  ⇒ {1, 2, 3, {}, 4, 5}
me:foo(1, 2, 3, 4, 5, 6)  ⇒ {1, 2, 3, {4}, 5, 6}
me:foo(1, 2, 3, 4, 5, 6, 7)  ⇒ {1, 2, 3, {4, 5}, 6, 7}
me:foo(1, 2, 3, 4, 5, 6, 7, 8) ⇒ {1, 2, 3, {4, 5, 6}, 7, 8}

Using scattering assignment, the example at the begining of this section could be rewritten more simply, reliably, and readably:

{first, second, ?third = 0} = args;

It is good MOO programming style to use a scattering assignment at the top of nearly every verb, since it shows so clearly just what kinds of arguments the verb expects.

4.1.10 Getting and Setting the Values of Properties

Usually, one can read the value of a property on an object with a simple expression:

expression.name Expression must return an object number; if not, E_TYPE is raised. If the object with that

number does not exist, E_INVIND is raised. Otherwise, if the object does not have a property with that name, then E_PROPNF is raised. Otherwise, if the named property is not readable by the owner of the current verb, then E_PERM is raised. Finally, assuming that none of these terrible things happens, the value of the named property on the given object is returned.

I said “usually” in the paragraph above because that simple expression only works if the name of the property obeys the same rules as for the names of variables (i.e., consists entirely of letters, digits, and underscores, and doesn’t begin with a digit). Property names are not restricted to this set, though. Also, it is sometimes useful to be able to figure out what property to read by some computation. For these more general uses, the following syntax is also allowed:

expression-1.(expression-2) As before, expression-1 must return an object number. Expression-2 must return a string, the

name of the property to be read; E_TYPE is raised otherwise. Using this syntax, any property can be read, regardless of its name.

Note that, as with almost everything in MOO, case is not significant in the names of properties. Thus, the following expressions are all equivalent:

     foo.bar
     foo.Bar
     foo.("bAr")

The LambdaCore database uses several properties on #0, the system object, for various special purposes. For example, the value of #0.room is the “generic room” object, #0.exit is the “generic exit” object, etc. This allows MOO programs to refer to these useful objects more easily (and more readably) than using their object numbers directly. To make this usage even easier and more readable, the expression

$name (where name obeys the rules for variable names) is an abbreviation for

#0.name Thus, for example, the value $nothing mentioned earlier is really #-1, the value of #0.nothing.

As with variables, one uses the assignment operator (‘=’) to change the value of a property. For example, the expression

     14 + (#27.foo = 17)

changes the value of the ‘foo’ property of the object numbered 27 to be 17 and then returns 31. Assignments to properties check that the owner of the current verb has write permission on the given property, raising E_PERM otherwise. Read permission is not required.

4.1.11 Calling Built-in Functions and Other Verbs

MOO provides a large number of useful functions for performing a wide variety of operations; a complete list, giving their names, arguments, and semantics, appears in a separate section later. As an example to give you the idea, there is a function named ‘length’ that returns the length of a given string or list.

The syntax of a call to a function is as follows:

name(expr-1, expr-2, …, expr-N) where name is the name of one of the built-in functions. The expressions between the parentheses, called arguments, are each evaluated in turn and then given to the named function to use in its appropriate way. Most functions require that a specific number of arguments be given; otherwise, E_ARGS is raised. Most also require that certain of the arguments have certain specified types (e.g., the length() function requires a list or a string as its argument); E_TYPE is raised if any argument has the wrong type.

As with list construction, the splicing operator ‘@’ can precede any argument expression. The value of such an expression must be a list; E_TYPE is raised otherwise. The elements of this list are passed as individual arguments, in place of the list as a whole.

Verbs can also call other verbs, usually using this syntax:

expr-0:name(expr-1, expr-2, …, expr-N) Expr-0 must return an object number; E_TYPE is raised otherwise. If the object with that number does not exist, E_INVIND is raised. If this task is too deeply nested in verbs calling verbs calling verbs, then E_MAXREC is raised; the default limit is 50 levels, but this can be changed from within the database; see the chapter on server assumptions about the database for details. If neither the object nor any of its ancestors defines a verb matching the given name, E_VERBNF is raised. Otherwise, if none of these nasty things happens, the named verb on the given object is called; the various built-in variables have the following initial values in the called verb:

this an object, the value of expr-0

verb a string, the name used in calling this verb

args a list, the values of expr-1, expr-2, etc.

caller an object, the value of this in the calling verb

player an object, the same value as it had initially in the calling verb or, if the calling verb is running with wizard permissions, the same as the current value in the calling verb.

All other built-in variables (argstr, dobj, etc.) are initialized with the same values they have in the calling verb.

As with the discussion of property references above, I said “usually” at the beginning of the previous paragraph because that syntax is only allowed when the name follows the rules for allowed variable names. Also as with property reference, there is a syntax allowing you to compute the name of the verb:

expr-0:(expr-00)(expr-1, expr-2, …, expr-N) The expression expr-00 must return a string; E_TYPE is raised otherwise.

The splicing operator (‘@’) can be used with verb-call arguments, too, just as with the arguments to built-in functions.

In many databases, a number of important verbs are defined on #0, the system object. As with the ‘$foo’ notation for properties on #0, the server defines a special syntax for calling verbs on #0:

$name(expr-1, expr-2, …, expr-N) (where name obeys the rules for variable names) is an abbreviation for

#0:name(expr-1, expr-2, …, expr-N)

4.1.12 Catching Errors in Expressions

It is often useful to be able to catch an error that an expression raises, to keep the error from aborting the whole task, and to keep on running as if the expression had returned some other value normally. The following expression accomplishes this:

` expr-1 ! codes => expr-2 ' Note: the open- and close-quotation marks in the previous line are really part of the syntax; you must actually type them as part of your MOO program for this kind of expression.

The codes part is either the keyword ANY or else a comma-separated list of expressions, just like an argument list. As in an argument list, the splicing operator (‘@’) can be used here. The => expr-2 part of the error-catching expression is optional.

First, the codes part is evaluated, yielding a list of error codes that should be caught if they’re raised; if codes is ANY, then it is equivalent to the list of all possible MOO values.

Next, expr-1 is evaluated. If it evaluates normally, without raising an error, then its value becomes the value of the entire error-catching expression. If evaluating expr-1 results in an error being raised, then call that error E. If E is in the list resulting from evaluating codes, then E is considered caught by this error-catching expression. In such a case, if expr-2 was given, it is evaluated to get the outcome of the entire error-catching expression; if expr-2 was omitted, then E becomes the value of the entire expression. If E is not in the list resulting from codes, then this expression does not catch the error at all and it continues to be raised, possibly to be caught by some piece of code either surrounding this expression or higher up on the verb-call stack. Here are some examples of the use of this kind of expression:

     `x + 1 ! E_TYPE => 0'

Returns x + 1 if x is an integer, returns 0 if x is not an integer, and raises E_VARNF if x doesn’t have a value.

     `x.y ! E_PROPNF, E_PERM => 17'

Returns x.y if that doesn’t cause an error, 17 if x doesn’t have a y property or that property isn’t readable, and raises some other kind of error (like E_INVIND) if x.y does.

     `1 / 0 ! ANY'

Returns E_DIV.

4.1.13 Parentheses and Operator Precedence

As shown in a few examples above, MOO allows you to use parentheses to make it clear how you intend for complex expressions to be grouped. For example, the expression

     3 * (4 + 5)

performs the addition of 4 and 5 before multiplying the result by 3.

If you leave out the parentheses, MOO will figure out how to group the expression according to certain rules. The first of these is that some operators have higher precedence than others; operators with higher precedence will more tightly bind to their operands than those with lower precedence. For example, multiplication has higher precedence than addition; thus, if the parentheses had been left out of the expression in the previous paragraph, MOO would have grouped it as follows:

     (3 * 4) + 5

The table below gives the relative precedence of all of the MOO operators; operators on higher lines in the table have higher precedence and those on the same line have identical precedence:

! - (without a left operand)

     ^
     *       /       %
     +       -
     ==      !=      <       <=      >       >=      in
     &&      ||

… ? … | … (the conditional expression)

     =

Thus, the horrendous expression

     x = a < b && c > d + e * f ? w in y | - q - r

would be grouped as follows:

     x = (((a < b) && (c > (d + (e * f)))) ? (w in y) | ((- q) - r))

It is best to keep expressions simpler than this and to use parentheses liberally to make your meaning clear to other humans.

4.2 MOO Language Statements

Statements are MOO constructs that, in contrast to expressions, perform some useful, nonvalue-producing operation. For example, there are several kinds of statements, called ‘looping constructs’, that repeatedly perform some set of operations. Fortunately, there are many fewer kinds of statements in MOO than there are kinds of expressions.

4.2.1 Errors While Executing Statements

Statements do not return values, but some kinds of statements can, under certain circumstances described below, generate errors. If such an error is generated in a verb whose ‘d’ (debug) bit is not set, then the error is ignored and the statement that generated it is simply skipped; execution proceeds with the next statement.

Note: this error-ignoring behavior is very error prone, since it affects all errors, including ones the programmer may not have anticipated. The ‘d’ bit exists only for historical reasons; it was once the only way for MOO programmers to catch and handle errors. The error-catching expression and the try-except statement are far better ways of accomplishing the same thing.

If the ‘d’ bit is set, as it usually is, then the error is raised and can be caught and handled either by code surrounding the expression in question or by verbs higher up on the chain of calls leading to the current verb. If the error is not caught, then the server aborts the entire task and, by default, prints a message to the current player. See the descriptions of the error-catching expression and the try-except statement for the details of how errors can be caught, and the chapter on server assumptions about the database for details on the handling of uncaught errors.

4.2.2 Simple Statements

The simplest kind of statement is the null statement, consisting of just a semicolon:

     ;

It doesn’t do anything at all, but it does it very quickly.

The next simplest statement is also one of the most common, the expression statement, consisting of any expression followed by a semicolon:

expression; The given expression is evaluated and the resulting value is ignored. Commonly-used kinds of expressions for such statements include assignments and verb calls. Of course, there’s no use for such a statement unless the evaluation of expression has some side-effect, such as changing the value of some variable or property, printing some text on someone’s screen, etc.

4.2.3 Statements for Testing Conditions

The ‘if’ statement allows you to decide whether or not to perform some statements based on the value of an arbitrary expression:

if (expression) statements

     endif

Expression is evaluated and, if it returns a true value, the statements are executed in order; otherwise, nothing more is done.

One frequently wants to perform one set of statements if some condition is true and some other set of statements otherwise. The optional ‘else’ phrase in an ‘if’ statement allows you to do this:

if (expression) statements-1

     else

statements-2

     endif

This statement is executed just like the previous one, except that statements-1 are executed if expression returns a true value and statements-2 are executed otherwise.

Sometimes, one needs to test several conditions in a kind of nested fashion:

if (expression-1) statements-1

     else

if (expression-2) statements-2

       else

if (expression-3) statements-3

         else

statements-4

         endif
       endif
     endif

Such code can easily become tedious to write and difficult to read. MOO provides a somewhat simpler notation for such cases: if (expression-1)

statements-1 elseif (expression-2)

statements-2 elseif (expression-3)

statements-3

else

statements-4

endif

Note that ‘elseif’ is written as a single word, without any spaces. This simpler version has the very same meaning as the original: evaluate expression-i for i equal to 1, 2, and 3, in turn, until one of them returns a true value; then execute the statements-i associated with that expression. If none of the expression-i return a true value, then execute statements-4.

Any number of ‘elseif’ phrases can appear, each having this form:

elseif (expression) statements

The complete syntax of the ‘if’ statement, therefore, is as follows:

if (expression)

statements zero-or-more-elseif-phrases an-optional-else-phrase

endif

4.2.4 Statements for Looping

MOO provides three different kinds of looping statements, allowing you to have a set of statements executed (1) once for each element of a given list, (2) once for each integer or object number in a given range, and (3) over and over until a given condition stops being true.

To perform some statements once for each element of a given list, use this syntax:

for variable in (expression)

statements

endfor

The expression is evaluated and should return a list; if it does not, E_TYPE is raised. The statements are then executed once for each element of that list in turn; each time, the given variable is assigned the value of the element in question. For example, consider the following statements:

odds = {1, 3, 5, 7, 9};
evens = {};
for n in (odds)
  evens = {@evens, n + 1};
endfor

The value of the variable ‘evens’ after executing these statements is the list

{2, 4, 6, 8, 10}

To perform a set of statements once for each integer or object number in a given range, use this syntax: for variable in [expression-1..expression-2]

statements

endfor

The two expressions are evaluated in turn and should either both return integers or both return object numbers; E_TYPE is raised otherwise. The statements are then executed once for each integer (or object number, as appropriate) greater than or equal to the value of expression-1 and less than or equal to the result of expression-2, in increasing order. Each time, the given variable is assigned the integer or object number in question. For example, consider the following statements:

evens = {};
for n in [1..5]
  evens = {@evens, 2 * n};
endfor

The value of the variable ‘evens’ after executing these statements is just as in the previous example: the list

{2, 4, 6, 8, 10}

The following loop over object numbers prints out the number and name of every valid object in the database:

for o in [#0..max_object()]
  if (valid(o))
    notify(player, tostr(o, ": ", o.name));
  endif
endfor

The final kind of loop in MOO executes a set of statements repeatedly as long as a given condition remains true:

while (expression)

statements

endwhile

The expression is evaluated and, if it returns a true value, the statements are executed; then, execution of the ‘while’ statement begins all over again with the evaluation of the expression. That is, execution alternates between evaluating the expression and executing the statements until the expression returns a false value. The following example code has precisely the same effect as the loop just shown above:

evens = {};
n = 1;
while (n <= 5)
  evens = {@evens, 2 * n};
  n = n + 1;
endwhile

Fine point: It is also possible to give a ‘name’ to a ‘while’ loop, using this syntax:

while name (expression)

statements

     endwhile

which has precisely the same effect as

while (name = expression) statements

          endwhile

This naming facility is only really useful in conjunction with the ‘break’ and ‘continue’ statements, described in the next section.

With each kind of loop, it is possible that the statements in the body of the loop will never be executed at all. For iteration over lists, this happens when the list returned by the expression is empty. For iteration on integers, it happens when expression-1 returns a larger integer than expression-2. Finally, for the ‘while’ loop, it happens if the expression returns a false value the very first time it is evaluated.

4.2.5 Terminating One or All Iterations of a Loop

Sometimes, it is useful to exit a loop before it finishes all of its iterations. For example, if the loop is used to search for a particular kind of element of a list, then it might make sense to stop looping as soon as the right kind of element is found, even if there are more elements yet to see. The ‘break’ statement is used for this purpose; it has the form

     break;

or

     break name;

Each ‘break’ statement indicates a specific surrounding loop; if name is not given, the statement refers to the innermost one. If it is given, name must be the name appearing right after the ‘for’ or ‘while’ keyword of the desired enclosing loop. When the ‘break’ statement is executed, the indicated loop is immediately terminated and executing continues just as if the loop had completed its iterations normally.

MOO also allows you to terminate just the current iteration of a loop, making it immediately go on to the next one, if any. The ‘continue’ statement does this; it has precisely the same forms as the ‘break’ statement:

     continue;

or

     continue name;

4.2.6 Returning a Value from a Verb

The MOO program in a verb is just a sequence of statements. Normally, when the verb is called, those statements are simply executed in order and then the integer 0 is returned as the value of the verb-call expression. Using the ‘return’ statement, one can change this behavior. The ‘return’ statement has one of the following two forms:

     return;

or

return expression; When it is executed, execution of the current verb is terminated immediately after evaluating the given expression, if any. The verb-call expression that started the execution of this verb then returns either the value of expression or the integer 0, if no expression was provided.

4.2.7 Handling Errors in Statements

Normally, whenever a piece of MOO code raises an error, the entire task is aborted and a message printed to the user. Often, such errors can be anticipated in advance by the programmer and code written to deal with them in a more graceful manner. The try-except statement allows you to do this; the syntax is as follows:

     try

statements-0 except variable-1 (codes-1)

statements-1 except variable-2 (codes-2)

statements-2 …

     endtry

where the variables may be omitted and each codes part is either the keyword ANY or else a commaseparated list of expressions, just like an argument list. As in an argument list, the splicing operator (‘@’) can be used here. There can be anywhere from 1 to 255 except clauses.

First, each codes part is evaluated, yielding a list of error codes that should be caught if they’re raised; if a codes is ANY, then it is equivalent to the list of all possible MOO values.

Next, statements-0 is executed; if it doesn’t raise an error, then that’s all that happens for the entire try-except statement. Otherwise, let E be the error it raises. From top to bottom, E is searched for in the lists resulting from the various codes parts; if it isn’t found in any of them, then it continues to be raised, possibly to be caught by some piece of code either surrounding this try-except statement or higher up on the verb-call stack.

If E is found first in codes-i, then variable-i (if provided) is assigned a value containing information about the error being raised and statements-i is executed. The value assigned to variable-i is a list of four elements:

{code, message, value, traceback} where code is E, the error being raised, message and value are as provided by the code that raised the error, and traceback is a list like that returned by the ‘callers()’ function, including line numbers. The traceback list contains entries for every verb from the one that raised the error through the one containing this try-except statement.

Unless otherwise mentioned, all of the built-in errors raised by expressions, statements, and functions provide tostr(code) as message and zero as value.

Here’s an example of the use of this kind of statement:

     try
       result = object:(command)(@arguments);
       player:tell("=> ", toliteral(result));
     except v (ANY)
       tb = v[4];
       if (length(tb) == 1)
         player:tell("** Illegal command: ", v[2]);
       else
         top = tb[1];
         tb[1..1] = {};
         player:tell(top[1], ":", top[2], ", line ", top[6], ":",

                     v[2]);
         for fr in (tb)
           player:tell("... called from ", fr[1], ":", fr[2],
                       ", line ", fr[6]);
         endfor
         player:tell("(End of traceback)");
       endif
     endtry

4.2.8 Cleaning Up After Errors

Whenever an error is raised, it is usually the case that at least some MOO code gets skipped over and never executed. Sometimes, it’s important that a piece of code always be executed, whether or not an error is raised. Use the try-finally statement for these cases; it has the following syntax:

     try

statements-1

     finally

statements-2

     endtry

First, statements-1 is executed; if it completes without raising an error, returning from this verb, or terminating the current iteration of a surrounding loop (we call these possibilities transferring control), then statements-2 is executed and that’s all that happens for the entire try-finally statement.

Otherwise, the process of transferring control is interrupted and statments-2 is executed. If statements-2 itself completes without transferring control, then the interrupted control transfer is resumed just where it left off. If statements-2 does transfer control, then the interrupted transfer is simply forgotten in favor of the new one.

In short, this statement ensures that statements-2 is executed after control leaves statements-1 for whatever reason; it can thus be used to make sure that some piece of cleanup code is run even if statements-1 doesn’t simply run normally to completion.

Here’s an example:

     try
       start = time();
       object:(command)(@arguments);
     finally
       end = time();
       this:charge_user_for_seconds(player, end - start);
     endtry

4.2.9 Executing Statements at a Later Time

It is sometimes useful to have some sequence of statements execute at a later time, without human intervention. For example, one might implement an object that, when thrown into the air, eventually falls back to the ground; the ‘throw’ verb on that object should arrange to print a message about the object landing on the ground, but the message shouldn’t be printed until some number of seconds have passed.

The ‘fork’ statement is intended for just such situations and has the following syntax: fork (expression) statements

     endfork

The ‘fork’ statement first executes the expression, which must return a integer; call that integer n. It then creates a new MOO task that will, after at least n seconds, execute the statements. When the new task begins, all variables will have the values they had at the time the ‘fork’ statement was executed. The task executing the ‘fork’ statement immediately continues execution. The concept of tasks is discussed in detail in the next section.

By default, there is no limit to the number of tasks any player may fork, but such a limit can be imposed from within the database. See the chapter on server assumptions about the database for details.

Occasionally, one would like to be able to kill a forked task before it even starts; for example, some player might have caught the object that was thrown into the air, so no message should be printed about it hitting the ground. If a variable name is given after the ‘fork’ keyword, like this:

fork name (expression)

statements

     endfork

then that variable is assigned the task ID of the newly-created task. The value of this variable is visible both to the task executing the fork statement and to the statements in the newly-created task. This ID can be passed to the kill_task() function to keep the task from running and will be the value of task_id() once the task begins execution.

4.3 MOO Tasks

A task is an execution of a MOO program. There are five kinds of tasks in LambdaMOO:

• Every time a player types a command, a task is created to execute that command; we call these command tasks.

• Whenever a player connects or disconnects from the MOO, the server starts a task to do whatever processing is necessary, such as printing out ‘Munchkin has connected’ to all of the players in the same room; these are called server tasks.

• The ‘fork’ statement in the programming language creates a task whose execution is delayed for at least some given number of seconds; these are forked tasks.

• The suspend() function suspends the execution of the current task. A snapshot is taken of whole state of the execution, and the execution will be resumed later. These are called suspended tasks.

• The read() function also suspends the execution of the current task, in this case waiting for the player to type a line of input. When the line is received, the task resumes with the read() function returning the input line as result. These are called reading tasks.

The last three kinds of tasks above are collectively known as queued tasks or background tasks, since they may not run immediately.

To prevent a maliciously- or incorrectly-written MOO program from running forever and monopolizing the server, limits are placed on the running time of every task. One limit is that no task is allowed to run longer than a certain number of seconds; command and server tasks get five seconds each while other tasks get only three seconds. This limit is, in practice, rarely reached. The reason is that there is also a limit on the number of operations a task may execute.

The server counts down ticks as any task executes. Roughly speaking, it counts one tick for every expression evaluation (other than variables and literals), one for every ‘if’, ‘fork’ or ‘return’ statement, and one for every iteration of a loop. If the count gets all the way down to zero, the task is immediately and unceremoniously aborted. By default, command and server tasks begin with an store of 30,000 ticks; this is enough for almost all normal uses. Forked, suspended, and reading tasks are allotted 15,000 ticks each.

These limits on seconds and ticks may be changed from within the database, as can the behavior of the server after it aborts a task for running out; see the chapter on server assumptions about the database for details.

Because queued tasks may exist for long periods of time before they begin execution, there are functions to list the ones that you own and to kill them before they execute. These functions, among others, are discussed in the following section.

4.4 Built-in Functions

There are a large number of built-in functions available for use by MOO programmers. Each one is discussed in detail in this section. The presentation is broken up into subsections by grouping together functions with similar or related uses.

For most functions, the expected types of the arguments are given; if the actual arguments are not of these types, E_TYPE is raised. Some arguments can be of any type at all; in such cases, no type specification is given for the argument. Also, for most functions, the type of the result of the function is given. Some functions do not return a useful result; in such cases, the specification ‘none’ is used. A few functions can potentially return any type of value at all; in such cases, the specification ‘value’ is used.

Most functions take a certain fixed number of required arguments and, in some cases, one or two optional arguments. If a function is called with too many or too few arguments, E_ARGS is raised.

Functions are always called by the program for some verb; that program is running with the permissions of some player, usually the owner of the verb in question (it is not always the owner, though; wizards can use set_task_perms() to change the permissions ‘on the fly’). In the function descriptions below, we refer to the player whose permissions are being used as the programmer.

Many built-in functions are described below as raising E_PERM unless the programmer meets certain specified criteria. It is possible to restrict use of any function, however, so that only wizards can use it; see the chapter on server assumptions about the database for details.

4.4.1 Object-Oriented Programming

One of the most important facilities in an object-oriented programming language is ability for a child object to make use of a parent’s implementation of some operation, even when the child provides its own definition for that operation. The pass() function provides this facility in MOO.

Function value pass (arg, …)

Often, it is useful for a child object to define a verb that augments the behavior of a verb on its parent object. For example, in the LambdaCore database, the root object (which is an ancestor of every other object) defines a verb called ‘description’ that simply returns the value of this.description; this verb is used by the implementation of the look command. In many cases, a programmer would like the description of some object to include some non-constant part; for example, a sentence about whether or not the object was ‘awake’ or ‘sleeping’. This sentence should be added onto the end of the normal description. The programmer would like to have a means of calling the normal description verb and then appending the sentence onto the end of that description. The function ‘pass()’ is for exactly such situations.

pass calls the verb with the same name as the current verb but as defined on the parent of the object that defines the current verb. The arguments given to pass are the ones given to the called verb and the returned value of the called verb is returned from the call to pass. The initial value of this in the called verb is the same as in the calling verb.

Thus, in the example above, the child-object’s description verb might have the following implementation:

     return pass() + "  It is " + (this.awake ? "awake." | "sleeping.");

That is, it calls its parent’s description verb and then appends to the result a sentence whose content is computed based on the value of a property on the object.

In almost all cases, you will want to call ‘pass()’ with the same arguments as were given to the current verb. This is easy to write in MOO; just call pass(@args).

4.4.2 Manipulating MOO Values

There are several functions for performing primitive operations on MOO values, and they can be cleanly split into two kinds: those that do various very general operations that apply to all types of values, and those that are specific to one particular type. There are so many operations concerned with objects that we do not list them in this section but rather give them their own section following this one.

4.4.2.1 General Operations Applicable to all Values

Function int typeof (value)

Takes any MOO value and returns an integer representing the type of value. The result is the same as the initial value of one of these built-in variables: INT, FLOAT, STR, LIST, OBJ, or ERR. Thus, one usually writes code like this:

     if (typeof(x) == LIST) . . .

and not like this:

     if (typeof(x) == 3) . . .

because the former is much more readable than the latter.

Function str tostr (value, …)

Converts all of the given MOO values into strings and returns the concatenation of the results.

tostr(17)  ⇒ "17"
tostr(1.0/3.0)  ⇒ "0.333333333333333"
tostr(#17)  ⇒ "#17"
tostr("foo")  ⇒ "foo"
tostr({1, 2})  ⇒ "{list}"
tostr(E_PERM)  ⇒ "Permission denied"
tostr("3 + 4 = ", 3 + 4) ⇒ "3 + 4 = 7"

Note that tostr() does not do a good job of converting lists into strings; all lists, including the empty list, are converted into the string "{list}". The function toliteral(), below, is better for this purpose.

Function str toliteral (value)

Returns a string containing a MOO literal expression that, when evaluated, would be equal to value.

toliteral(17)  ⇒ "17"
toliteral(1.0/3.0)  ⇒ "0.333333333333333"
toliteral(#17)  ⇒ "#17"
toliteral("foo")  ⇒ "\"foo\""
toliteral({1, 2})  ⇒ "{1, 2}"
toliteral(E_PERM)  ⇒ "E_PERM"

Function int toint (value) Function int tonum (value)

Converts the given MOO value into an integer and returns that integer. Floating-point numbers are rounded toward zero, truncating their fractional parts. Object numbers are converted into the equivalent integers. Strings are parsed as the decimal encoding of a real number which is then converted to an integer. Errors are converted into integers obeying the same ordering (with respect to <= as the errors themselves. Toint() raises E_TYPE if value is a list. If value is a string but the string does not contain a syntactically-correct number, then toint() returns 0.

toint(34.7)  ⇒ 34
toint(-34.7)  ⇒ -34
toint(#34)  ⇒ 34
toint("34")  ⇒ 34
toint("34.7")  ⇒ 34
toint(" - 34  ") ⇒ -34
toint(E_TYPE)  ⇒ 1

Function obj toobj (value)

Converts the given MOO value into an object number and returns that object number. The conversions are very similar to those for toint() except that for strings, the number may be preceded by ‘#’.

toobj("34")  ⇒ #34
toobj("#34")  ⇒ #34
toobj("foo")  ⇒ #0
toobj({1, 2})      error  E_TYPE

Function float tofloat (value)

Converts the given MOO value into a floating-point number and returns that number. Integers and object numbers are converted into the corresponding integral floatingpoint numbers. Strings are parsed as the decimal encoding of a real number which is then represented as closely as possible as a floating-point number. Errors are first converted to integers as in toint() and then converted as integers are. Tofloat() raises E_TYPE if value is a list. If value is a string but the string does not contain a syntactically-correct number, then tofloat() returns 0.

     tofloat(34)  ⇒ 34.0
     tofloat(#34)  ⇒ 34.0
     tofloat("34")  ⇒ 34.0
     tofloat("34.7")  ⇒ 34.7
     tofloat(E_TYPE)  ⇒ 1.0

Function int equal (value1, value2)

Returns true if value1 is completely indistinguishable from value2. This is much the same operation as “value1 == value2” except that, unlike ==, the equal() function does not treat upper- and lower-case characters in strings as equal.

     "Foo" == "foo"  ⇒ 1
     equal("Foo", "foo")  ⇒ 0
     equal("Foo", "Foo")  ⇒ 1

Function int value bytes (value)

Returns the number of bytes of the server’s memory required to store the given value.

Function str value hash (value)

Returns the same string as string_hash(toliteral(value)); see the description of

string_hash() for details.

4.4.2.2 Operations on Numbers

Function int random ([int mod])

mod must be a positive integer; otherwise, E_INVARG is raised. An integer is chosen randomly from the range [1..mod] and returned. If mod is not provided, it defaults to the largest MOO integer, 2147483647.

Function num min (num x, …) Function num max (num x, …)

These two functions return the smallest or largest of their arguments, respectively. All of the arguments must be numbers of the same kind (i.e., either integer or floatingpoint); otherwise E_TYPE is raised.

Function num abs (num x)

Returns the absolute value of x. If x is negative, then the result is -x; otherwise, the result is x. The number x can be either integer or floating-point; the result is of the same kind.

Function str floatstr(float x, int precision [, scientific])

Converts x into a string with more control than provided by either tostr() or toliteral(). Precision is the number of digits to appear to the right of the decimal point, capped at 4 more than the maximum available precision, a total of 19 on most machines; this makes it possible to avoid rounding errors if the resulting string is subsequently read back as a floating-point value. If scientific is false or not provided, the result is a string in the form "MMMMMMM.DDDDDD", preceded by a minus sign if and only if x is negative. If scientific is provided and true, the result is a string in the form "M.DDDDDDe+EEE", again preceded by a minus sign if and only if x is negative.

Function float sqrt (float x)

Returns the square root of x. Raises E_INVARG if x is negative.

Function float sin (float x) Function float cos (float x) Function float tan (float x)

Returns the sine, cosine, or tangent of x, respectively.

Function float asin (float x) Function float acos (float x)

Returns the arc-sine or arc-cosine (inverse sine or cosine) of x, in the range [- pi/2..pi/2] or [0..pi], respectively. Raises E_INVARG if x is outside the range

[-1.0..1.0].

Function float atan (float y [, float x])

Returns the arc-tangent (inverse tangent) of y in the range [-pi/2..pi/2] if x is not provided, or of y/x in the range [-pi..pi] if x is provided.

Function float sinh (float x) Function float cosh (float x) Function float tanh (float x)

Returns the hyperbolic sine, cosine, or tangent of x, respectively.

Function float exp (float x)

Returns e raised to the power of x.

Function float log (float x) Function float log10 (float x)

Returns the natural or base 10 logarithm of x. Raises E_INVARG if x is not positive.

Function float ceil (float x)

Returns the smallest integer not less than x, as a floating-point number.

Function float floor (float x)

Returns the largest integer not greater than x, as a floating-point number.

Function float trunc (float x)

Returns the integer obtained by truncating x at the decimal point, as a floating-point number. For negative x, this is equivalent to ceil(); otherwise it is equivalent to

floor().

4.4.2.3 Operations on Strings

Function int length (str string)

Returns the number of characters in string. It is also permissible to pass a list to length(); see the description in the next section.

     length("foo") ⇒ 3
     length("")  ⇒ 0

Function str strsub (str subject, str what, str with [, case-matters])

Replaces all occurrences in subject of what with with, performing string substitution. The occurrences are found from left to right and all substitutions happen simultaneously. By default, occurrences of what are searched for while ignoring the upper/lower case distinction. If case-matters is provided and true, then case is treated as significant in all comparisons.

     strsub("%n is a fink.", "%n", "Fred") ⇒ "Fred is a fink."
     strsub("foobar", "OB", "b")  ⇒ "fobar"
     strsub("foobar", "OB", "b", 1)  ⇒ "foobar"

Function int index (str str1, str str2 [, case-matters]) Function int rindex (str str1, str str2 [, case-matters])

The function index() (rindex()) returns the index of the first character of the first (last) occurrence of str2 in str1, or zero if str2 does not occur in str1 at all. By default the search for an occurrence of str2 is done while ignoring the upper/lower case distinction. If case-matters is provided and true, then case is treated as significant in all comparisons.

     index("foobar", "o")  ⇒ 2
     rindex("foobar", "o")  ⇒ 3
     index("foobar", "x")  ⇒ 0
     index("foobar", "oba")  ⇒ 3
     index("Foobar", "foo", 1) ⇒ 0

Function int strcmp (str str1, str str2)

Performs a case-sensitive comparison of the two argument strings. If str1 is lexicographically less than str2, the strcmp() returns a negative integer. If the two strings are identical, strcmp() returns zero. Otherwise, strcmp() returns a positive integer. The ASCII character ordering is used for the comparison.

Function list decode binary (str bin-string [, fully])

Returns a list of strings and/or integers representing the bytes in the binary string bin string in order. If fully is false or omitted, the list contains an integer only for each non-printing, non-space byte; all other characters are grouped into the longest possible contiguous substrings. If fully is provided and true, the list contains only integers, one for each byte represented in bin string. Raises E_INVARG if bin string is not a properlyformed binary string. (See the early section on MOO value types for a full description of binary strings.)

     decode_binary("foo")  ⇒ {"foo"}
     decode_binary("~~foo")  ⇒ {"~foo"}
     decode_binary("foo~0D~0A")  ⇒ {"foo", 13, 10}

     decode_binary("foo~0Abar~0Abaz") ⇒ {"foo", 10, "bar", 10, "baz"}
     decode_binary("foo~0D~0A", 1)  ⇒ {102, 111, 111, 13, 10}

Function str encode binary (arg, …)

Each argument must be an integer between 0 and 255, a string, or a list containing only legal arguments for this function. This function translates each integer and string in turn into its binary string equivalent, returning the concatenation of all these substrings into a single binary string. (See the early section on MOO value types for a full description of binary strings.)

     encode_binary("~foo")  ⇒ "~7Efoo"
     encode_binary({"foo", 10}, {"bar", 13}) ⇒ "foo~0Abar~0D"
     encode_binary("foo", 10, "bar", 13)  ⇒ "foo~0Abar~0D"

Function list match (str subject, str pattern [, case-matters]) Function list rmatch (str subject, str pattern [, case-matters])

The function match() (rmatch()) searches for the first (last) occurrence of the regular expression pattern in the string subject. If pattern is syntactically malformed, then E_INVARG is raised. The process of matching can in some cases consume a great deal of memory in the server; should this memory consumption become excessive, then the matching process is aborted and E_QUOTA is raised. If no match is found, the empty list is returned; otherwise, these functions return a list containing information about the match (see below). By default, the search ignores upper-/lower-case distinctions. If case-matters is provided and true, then case is treated as significant in all comparisons. The list that match() (rmatch()) returns contains the details about the match made. The list is in the form:

{start, end, replacements, subject} where start is the index in subject of the beginning of the match, end is the index of the end of the match, replacements is a list described below, and subject is the same string that was given as the first argument to the match() or rmatch(). The replacements list is always nine items long, each item itself being a list of two integers, the start and end indices in string matched by some parenthesized sub-pattern of pattern. The first item in replacements carries the indices for the first parenthesized sub-pattern, the second item carries those for the second sub-pattern, and so on. If there are fewer than nine parenthesized sub-patterns in pattern, or if some sub-pattern was not used in the match, then the corresponding item in replacements is the list {0, -1}. See the discussion of ‘%)’, below, for more information on parenthesized subpatterns.

     match("foo", "^f*o$")  ⇒ {}
     match("foo", "^fo*$")  ⇒ {1, 3, {{0, -1}, . . .}, "foo"}
     match("foobar", "o*b")  ⇒ {2, 4, {{0, -1}, . . .}, "foobar"}
     rmatch("foobar", "o*b")  ⇒ {4, 4, {{0, -1}, . . .}, "foobar"}
     match("foobar", "f%(o*%)b")
             ⇒ {1, 4, {{2, 3}, {0, -1}, . . .}, "foobar"}

Regular expression matching allows you to test whether a string fits into a specific syntactic shape. You can also search a string for a substring that fits a pattern.

A regular expression describes a set of strings. The simplest case is one that describes a particular string; for example, the string ‘foo’ when regarded as a regular expression matches ‘foo’ and nothing else. Nontrivial regular expressions use certain special constructs so that they can match more than one string. For example, the regular expression ‘foo%|bar’ matches either the string ‘foo’ or the string ‘bar’; the regular expression ‘c[ad]*r’ matches any of the strings ‘cr’, ‘car’, ‘cdr’, ‘caar’, ‘cadddar’ and all other such strings with any number of ‘a’’s and ‘d’’s. Regular expressions have a syntax in which a few characters are special constructs and the rest are ordinary. An ordinary character is a simple regular expression that matches that character and nothing else. The special characters are ‘$’, ‘^’, ‘.’, ‘*’, ‘+’, ‘?’, ‘[’, ‘]’ and ‘%’. Any other character appearing in a regular expression is ordinary, unless a ‘%’ precedes it. For example, ‘f’ is not a special character, so it is ordinary, and therefore ‘f’ is a regular expression that matches the string ‘f’ and no other string. (It does not, for example, match the string ‘ff’.) Likewise, ‘o’ is a regular expression that matches only ‘o’. Any two regular expressions a and b can be concatenated. The result is a regular expression which matches a string if a matches some amount of the beginning of that string and b matches the rest of the string. As a simple example, we can concatenate the regular expressions ‘f’ and ‘o’ to get the regular expression ‘fo’, which matches only the string ‘fo’. Still trivial. The following are the characters and character sequences that have special meaning within regular expressions. Any character not mentioned here is not special; it stands for exactly itself for the purposes of searching and matching.

‘.’ is a special character that matches any single character. Using concatenation, we can make regular expressions like ‘a.b’, which matches any three-character string that begins with ‘a’ and ends with ‘b’.

‘*’ is not a construct by itself; it is a suffix that means that the preceding regular expression is to be repeated as many times as possible. In ‘fo*’, the ‘*’ applies to the ‘o’, so ‘fo*’ matches ‘f’ followed by any number of ‘o’’s. The case of zero ‘o’’s is allowed: ‘fo*’ does match ‘f’. ‘*’ always applies to the smallest possible preceding expression. Thus, ‘fo*’ has a repeating ‘o’, not a repeating ‘fo’. The matcher processes a ‘*’ construct by matching, immediately, as many repetitions as can be found. Then it continues with the rest of the pattern. If that fails, it backtracks, discarding some of the matches of the ‘*’’d construct in case that makes it possible to match the rest of the pattern. For example, matching ‘c[ad]*ar’ against the string ‘caddaar’, the ‘[ad]*’ first matches ‘addaa’, but this does not allow the next ‘a’ in the pattern to match. So the last of the matches of ‘[ad]’ is undone and the following ‘a’ is tried again. Now it succeeds.

‘+’ ‘+’ is like ‘*’ except that at least one match for the preceding pattern is required for ‘+’. Thus, ‘c[ad]+r’ does not match ‘cr’ but does match anything else that ‘c[ad]*r’ would match.

‘?’ ‘?’ is like ‘*’ except that it allows either zero or one match for the preceding pattern. Thus, ‘c[ad]?r’ matches ‘cr’ or ‘car’ or ‘cdr’, and nothing else.

‘[ …]’ ‘[’ begins a character set, which is terminated by a ‘]’. In the simplest case, the characters between the two brackets form the set. Thus, ‘[ad]’ matches either ‘a’ or ‘d’, and ‘[ad]*’ matches any string of ‘a’’s and ‘d’’s (including the empty string), from which it follows that ‘c[ad]*r’ matches ‘car’, etc. Character ranges can also be included in a character set, by writing two characters with a ‘-’ between them. Thus, ‘[a-z]’ matches any lower-case letter. Ranges may be intermixed freely with individual characters, as in ‘[a-z$%.]’, which matches any lower case letter or ‘$’, ‘%’ or period. Note that the usual special characters are not special any more inside a character set. A completely different set of special characters exists inside character sets: ‘]’, ‘-’ and ‘^’. To include a ‘]’ in a character set, you must make it the first character. For example, ‘[]a]’ matches ‘]’ or ‘a’. To include a ‘-’, you must use it in a context where it cannot possibly indicate a range: that is, as the first character, or immediately after a range.

‘[^…]’ ‘[^’ begins a complement character set, which matches any character except the ones specified. Thus, ‘[^a-z0-9A-Z]’ matches all characters except letters and digits. ‘^’ is not special in a character set unless it is the first character. The character following the ‘^’ is treated as if it were first (it may be a ‘-’ or a ‘]’).

‘^’ is a special character that matches the empty string – but only if at the beginning of the string being matched. Otherwise it fails to match anything. Thus, ‘^foo’ matches a ‘foo’ which occurs at the beginning of the string.

‘$’ is similar to ‘^’ but matches only at the end of the string. Thus, ‘xx*$’ matches a string of one or more ‘x’’s at the end of the string.

‘%’ has two functions: it quotes the above special characters (including ‘%’), and it introduces additional special constructs. Because ‘%’ quotes special characters, ‘%$’ is a regular expression that matches only ‘$’, and ‘%[’ is a regular expression that matches only ‘[’, and so on. For the most part, ‘%’ followed by any character matches only that character. However, there are several exceptions: characters that, when preceded by ‘%’, are special constructs. Such characters are always ordinary when encountered on their own. No new special characters will ever be defined. All extensions to the regular expression syntax are made by defining new two-character constructs that begin with ‘%’.

‘%|’ specifies an alternative. Two regular expressions a and b with ‘%|’ in between form an expression that matches anything that either a or b will match.

Thus, ‘foo%|bar’ matches either ‘foo’ or ‘bar’ but no other string. ‘%|’ applies to the largest possible surrounding expressions. Only a surrounding ‘%( … %)’ grouping can limit the grouping power of ‘%|’. Full backtracking capability exists for when multiple ‘%|’’s are used.

‘%( … %)’ is a grouping construct that serves three purposes:

1. To enclose a set of ‘%|’ alternatives for other operations. Thus, ‘%(foo%|bar%)x’ matches either ‘foox’ or ‘barx’.

2. To enclose a complicated expression for a following ‘*’, ‘+’, or ‘?’ to operate on. Thus, ‘ba%(na%)*’ matches ‘bananana’, etc., with any number of ‘na’’s, including none.

3. To mark a matched substring for future reference.

This last application is not a consequence of the idea of a parenthetical grouping; it is a separate feature that happens to be assigned as a second meaning to the same ‘%( … %)’ construct because there is no conflict in practice between the two meanings. Here is an explanation of this feature:

‘%digit’ After the end of a ‘%( … %)’ construct, the matcher remembers the beginning and end of the text matched by that construct. Then, later on in the regular expression, you can use ‘%’ followed by digit to mean “match the same text matched by the digit’th ‘%( … %)’ construct in the pattern.” The ‘%( … %)’ constructs are numbered in the order that their ‘%(’’s appear in the pattern. The strings matching the first nine ‘%( … %)’ constructs appearing in a regular expression are assigned numbers 1 through 9 in order of their beginnings. ‘%1’ through ‘%9’ may be used to refer to the text matched by the corresponding ‘%( … %)’ construct. For example, ‘%(.*%)%1’ matches any string that is composed of two identical halves. The ‘%(.*%)’ matches the first half, which may be anything, but the ‘%1’ that follows must match the same exact text.

‘%b’ matches the empty string, but only if it is at the beginning or end of a word. Thus, ‘%bfoo%b’ matches any occurrence of ‘foo’ as a separate word. ‘%bball%(s%|%)%b’ matches ‘ball’ or ‘balls’ as a separate word. For the purposes of this construct and the five that follow, a word is defined to be a sequence of letters and/or digits.

‘%B’ matches the empty string, provided it is not at the beginning or end of a word.

‘%<’ matches the empty string, but only if it is at the beginning of a word.

‘%>’ matches the empty string, but only if it is at the end of a word.

‘%w’ matches any word-constituent character (i.e., any letter or digit).

‘%W’ matches any character that is not a word constituent.

Function str substitute (str template, list subs)

Performs a standard set of substitutions on the string template, using the information contained in subs, returning the resulting, transformed template. Subs should be a list like those returned by match() or rmatch() when the match succeeds; otherwise, E_INVARG is raised.

In template, the strings ‘%1’ through ‘%9’ will be replaced by the text matched by the first through ninth parenthesized sub-patterns when match() or rmatch() was called. The string ‘%0’ in template will be replaced by the text matched by the pattern as a whole when match() or rmatch() was called. The string ‘%%’ will be replaced by a single ‘%’ sign. If ‘%’ appears in template followed by any other character, E_INVARG will be raised.

     subs = match("*** Welcome to LambdaMOO!!!", "%(%w*%) to %(%w*%)");
     substitute("I thank you for your %1 here in %2.", subs)
             ⇒ "I thank you for your Welcome here in LambdaMOO."

Function str crypt (str text [, str salt])

Encrypts the given text using the standard UNIX encryption method. If provided, salt should be a string at least two characters long, the first two characters of which will be used as the extra encryption “salt” in the algorithm. If salt is not provided, a random pair of characters is used. In any case, the salt used is also returned as the first two characters of the resulting encrypted string.

Aside from the possibly-random selection of the salt, the encryption algorithm is entirely deterministic. In particular, you can test whether or not a given string is the same as the one used to produce a given piece of encrypted text; simply extract the first two characters of the encrypted text and pass the candidate string and those two characters to crypt(). If the result is identical to the given encrypted text, then you’ve got a match.

     crypt("foobar")  ⇒ "J3fSFQfgkp26w"
     crypt("foobar", "J3") ⇒ "J3fSFQfgkp26w"
     crypt("mumble", "J3") ⇒ "J3D0.dh.jjmWQ"
     crypt("foobar", "J4") ⇒ "J4AcPxOJ4ncq2"

Function str string hash (str text)

Function str binary hash (str bin-string)

Returns a 32-character hexadecimal string encoding the result of applying the MD5 cryptographically secure hash function to the contents of the string text or the binary string bin-string. MD5, like other such functions, has the property that, if

     string_hash(x) == string_hash(y)

then, almost certainly,

     equal(x, y)

This can be useful, for example, in certain networking applications: after sending a large piece of text across a connection, also send the result of applying string_hash() to the text; if the destination site also applies string_hash() to the text and gets the same result, you can be quite confident that the large text has arrived unchanged.

4.4.2.4 Operations on Lists

Function int length (list list)

Returns the number of elements in list. It is also permissible to pass a string to length(); see the description in the previous section.

     length({1, 2, 3}) ⇒ 3

     length({})  ⇒ 0

Function int is member (value, list list)

Returns true if there is an element of list that is completely indistinguishable from value. This is much the same operation as “value in list” except that, unlike in, the is_member() function does not treat upper- and lower-case characters in strings as equal.

     "Foo" in {1, "foo", #24}  ⇒ 2
     is_member("Foo", {1, "foo", #24}) ⇒ 0
     is_member("Foo", {1, "Foo", #24}) ⇒ 2

Function list listinsert (list list, value [, int index]) Function list listappend (list list, value [, int index])

These functions return a copy of list with value added as a new element. listinsert() and listappend() add value before and after (respectively) the existing element with the given index, if provided. The following three expressions always have the same value:

listinsert(list, element, index)

     listappend(list, element, index - 1)

{@list[1..index - 1], element, @list[index..length(list)]} If index is not provided, then listappend() adds the value at the end of the list and listinsert() adds it at the beginning; this usage is discouraged, however, since the same intent can be more clearly expressed using the list-construction expression, as shown in the examples below.

     x = {1, 2, 3};
     listappend(x, 4, 2) ⇒ {1, 2, 4, 3}
     listinsert(x, 4, 2) ⇒ {1, 4, 2, 3}
     listappend(x, 4)  ⇒ {1, 2, 3, 4}
     listinsert(x, 4)  ⇒ {4, 1, 2, 3}
     {@x, 4}  ⇒ {1, 2, 3, 4}
     {4, @x}  ⇒ {4, 1, 2, 3}

Function list listdelete (list list, int index)

Returns a copy of list with the indexth element removed. If index is not in the range

[1..length(list)], then E_RANGE is raised.
     x = {"foo", "bar", "baz"};
     listdelete(x, 2) ⇒ {"foo", "baz"}

Function list listset (list list, value, int index)

Returns a copy of list with the indexth element replaced by value. If index is not in the range [1..length(list)], then E_RANGE is raised.

     x = {"foo", "bar", "baz"};
     listset(x, "mumble", 2) ⇒ {"foo", "mumble", "baz"}

This function exists primarily for historical reasons; it was used heavily before the server supported indexed assignments like x[i] = v. New code should always use indexed assignment instead of ‘listset()’ wherever possible.

Function list setadd (list list, value) Function list setremove (list list, value)

Returns a copy of list with the given value added or removed, as appropriate. setadd() only adds value if it is not already an element of list; list is thus treated as a mathematical set. value is added at the end of the resulting list, if at all. Similarly, setremove() returns a list identical to list if value is not an element. If value appears more than once in list, only the first occurrence is removed in the returned copy.

     setadd({1, 2, 3}, 3)  ⇒ {1, 2, 3}
     setadd({1, 2, 3}, 4)  ⇒ {1, 2, 3, 4}
     setremove({1, 2, 3}, 3)  ⇒ {1, 2}
     setremove({1, 2, 3}, 4)  ⇒ {1, 2, 3}
     setremove({1, 2, 3, 2}, 2) ⇒ {1, 3, 2}

4.4.3 Manipulating Objects

Objects are, of course, the main focus of most MOO programming and, largely due to that, there are a lot of built-in functions for manipulating them.

4.4.3.1 Fundamental Operations on Objects

Function obj create (obj parent [, obj owner])

Creates and returns a new object whose parent is parent and whose owner is as described below. Either the given parent object must be #-1 or valid and fertile (i.e., its ‘f’ bit must be set) or else the programmer must own parent or be a wizard; otherwise E_PERM is raised. E_PERM is also raised if owner is provided and not the same as the programmer, unless the programmer is a wizard. After the new object is created, its initialize verb, if any, is called with no arguments. The new object is assigned the least non-negative object number that has not yet been used for a created object. Note that no object number is ever reused, even if the object with that number is recycled. The owner of the new object is either the programmer (if owner is not provided), the new object itself (if owner was given as #-1), or owner (otherwise). The other built-in properties of the new object are initialized as follows:

     name         ""
     location     #-1
     contents     {}
     programmer   0
     wizard       0
     r            0

     w            0
     f            0

The function ‘is_player()’ returns false for newly created objects. In addition, the new object inherits all of the other properties on parent. These properties have the same permission bits as on parent. If the ‘c’ permissions bit is set, then the owner of the property on the new object is the same as the owner of the new object itself; otherwise, the owner of the property on the new object is the same as that on parent. The initial value of every inherited property is clear; see the description of the built-in function clear_property() for details. If the intended owner of the new object has a property named ‘ownership_quota’ and the value of that property is an integer, then create() treats that value as a quota. If the quota is less than or equal to zero, then the quota is considered to be exhausted and create() raises E_QUOTA instead of creating an object. Otherwise, the quota is decremented and stored back into the ‘ownership_quota’ property as a part of the creation of the new object.

Function none chparent (obj object, obj new-parent)

Changes the parent of object to be new-parent. If object is not valid, or if new-parent is neither valid nor equal to #-1, then E_INVARG is raised. If the programmer is neither a wizard or the owner of object, or if new-parent is not fertile (i.e., its ‘f’ bit is not set) and the programmer is neither the owner of new-parent nor a wizard, then E_PERM is raised. If new-parent is equal to object or one of its current ancestors, E_RECMOVE is raised. If object or one of its descendants defines a property with the same name as one defined either on new-parent or on one of its ancestors, then E_INVARG is raised. Changing an object’s parent can have the effect of removing some properties from and adding some other properties to that object and all of its descendants (i.e., its children and its children’s children, etc.). Let common be the nearest ancestor that object and new-parent have in common before the parent of object is changed. Then all properties defined by ancestors of object under common (that is, those ancestors of object that are in turn descendants of common) are removed from object and all of its descendants. All properties defined by new-parent or its ancestors under common are added to object and all of its descendants. As with create(), the newly-added properties are given the same permission bits as they have on new-parent, the owner of each added property is either the owner of the object it’s added to (if the ‘c’ permissions bit is set) or the owner of that property on new-parent, and the value of each added property is clear; see the description of the built-in function clear_property() for details. All properties that are not removed or added in the reparenting process are completely unchanged. If new-parent is equal to #-1, then object is given no parent at all; it becomes a new root of the parent/child hierarchy. In this case, all formerly inherited properties on object are simply removed.

Function int valid (obj object)

Returns a non-zero integer (i.e., a true value) if object is a valid object (one that has been created and not yet recycled) and zero (i.e., a false value) otherwise.

     valid(#0)  ⇒ 1

     valid(#-1) ⇒ 0

Function obj parent (obj object)

Function list children (obj object)

These functions return the parent and a list of the children of object, respectively. If object is not valid, then E_INVARG is raised.

Function none recycle (obj object)

The given object is destroyed, irrevocably. The programmer must either own object or be a wizard; otherwise, E_PERM is raised. If object is not valid, then E_INVARG is raised. The children of object are reparented to the parent of object. Before object is recycled, each object in its contents is moved to #-1 (implying a call to object’s exitfunc verb, if any) and then object’s ‘recycle’ verb, if any, is called with no arguments. After object is recycled, if the owner of the former object has a property named ‘ownership_quota’ and the value of that property is a integer, then recycle() treats that value as a quota and increments it by one, storing the result back into the

‘ownership_quota’ property.

Function int object bytes (obj object)

Returns the number of bytes of the server’s memory required to store the given object, including the space used by the values of all of its non-clear properties and by the verbs and properties defined directly on the object. Raised E_INVARG if object is not a valid object and E_PERM if the programmer is not a wizard.

Function obj max object ()

Returns the largest object number yet assigned to a created object. Note that the object with this number may no longer exist; it may have been recycled. The next object created will be assigned the object number one larger than the value of max_

object().

4.4.3.2 Object Movement

Function none move (obj what, obj where)

Changes what’s location to be where. This is a complex process because a number of permissions checks and notifications must be performed. The actual movement takes place as described in the following paragraphs. what should be a valid object and where should be either a valid object or #-1 (denoting a location of ‘nowhere’); otherwise E_INVARG is raised. The programmer must be either the owner of what or a wizard; otherwise, E_PERM is raised. If where is a valid object, then the verb-call

     where:accept(what)

is performed before any movement takes place. If the verb returns a false value and the programmer is not a wizard, then where is considered to have refused entrance to what; move() raises E_NACC. If where does not define an accept verb, then it is treated as if it defined one that always returned false. If moving what into where would create a loop in the containment hierarchy (i.e., what would contain itself, even indirectly), then E_RECMOVE is raised instead.

The ‘location’ property of what is changed to be where, and the ‘contents’ properties of the old and new locations are modified appropriately. Let old-where be the location of what before it was moved. If old-where is a valid object, then the verb-call

old-where:exitfunc(what) is performed and its result is ignored; it is not an error if old-where does not define a verb named ‘exitfunc’. Finally, if where and what are still valid objects, and where is still the location of what, then the verb-call

     where:enterfunc(what)

is performed and its result is ignored; again, it is not an error if where does not define a verb named ‘enterfunc’.

4.4.3.3 Operations on Properties

Function list properties (obj object)

Returns a list of the names of the properties defined directly on the given object, not inherited from its parent. If object is not valid, then E_INVARG is raised. If the programmer does not have read permission on object, then E_PERM is raised.

Function list property info (obj object, str prop-name) Function none set property info (obj object, str prop-name, list info)

These two functions get and set (respectively) the owner and permission bits for the property named prop-name on the given object. If object is not valid, then E_INVARG is raised. If object has no non-built-in property named prop-name, then E_PROPNF is raised. If the programmer does not have read (write) permission on the property

in question, then property_info() (set_property_info()) raises E_PERM. Property

info has the following form:

{owner, perms [, new-name]} where owner is an object, perms is a string containing only characters from the set ‘r’, ‘w’, and ‘c’, and new-name is a string; new-name is never part of the value returned by property_info(), but it may optionally be given as part of the value provided to set_property_info(). This list is the kind of value returned by property_info() and expected as the third argument to set_property_info(); the latter function raises E_INVARG if owner is not valid, if perms contains any illegal characters, or, when newname is given, if prop-name is not defined directly on object or new-name names an existing property defined on object or any of its ancestors or descendants.

Function none add property (obj object, str prop-name, value, list info)

Defines a new property on the given object, inherited by all of its descendants; the property is named prop-name, its initial value is value, and its owner and initial permission bits are given by info in the same format as is returned by property_info(), described above. If object is not valid or info does not specify a valid owner and well-formed permission bits or object or its ancestors or descendants already defines a property named prop-name, then E_INVARG is raised. If the programmer does not have write permission on object or if the owner specified by info is not the programmer and the programmer is not a wizard, then E_PERM is raised.

Function none delete property (obj object, str prop-name)

Removes the property named prop-name from the given object and all of its descendants. If object is not valid, then E_INVARG is raised. If the programmer does not have write permission on object, then E_PERM is raised. If object does not directly define a property named prop-name (as opposed to inheriting one from its parent), then E_PROPNF is raised.

Function int is clear property (obj object, str prop-name) Function none clear property (obj object, str prop-name)

These two functions test for clear and set to clear, respectively, the property named prop-name on the given object. If object is not valid, then E_INVARG is raised. If object has no non-built-in property named prop-name, then E_PROPNF is raised. If the programmer does not have read (write) permission on the property in question, then

is_clear_property() (clear_property()) raises E_PERM. If a property is clear, then

when the value of that property is queried the value of the parent’s property of the same name is returned. If the parent’s property is clear, then the parent’s parent’s value is examined, and so on. If object is the definer of the property prop-name, as opposed to an inheritor of the property, then clear_property() raises E_INVARG.

4.4.3.4 Operations on Verbs

Function list verbs (obj object)

Returns a list of the names of the verbs defined directly on the given object, not inherited from its parent. If object is not valid, then E_INVARG is raised. If the programmer does not have read permission on object, then E_PERM is raised.

Most of the remaining operations on verbs accept a string containing the verb’s name to identify the verb in question. Because verbs can have multiple names and because an object can have multiple verbs with the same name, this practice can lead to difficulties. To most unambiguously refer to a particular verb, one can instead use a positive integer, the index of the verb in the list returned by verbs(), described above.

For example, suppose that verbs(#34) returns this list:

{"foo", "bar", "baz", "foo"}

Object #34 has two verbs named ‘foo’ defined on it (this may not be an error, if the two verbs have different command syntaxes). To refer unambiguously to the first one in the list, one uses the integer 1; to refer to the other one, one uses 4.

In the function descriptions below, an argument named verb-desc is either a string containing the name of a verb or else a positive integer giving the index of that verb in its defining object’s

verbs() list.

For historical reasons, there is also a second, inferior mechanism for referring to verbs with numbers, but its use is strongly discouraged. If the property $server_

options.support_numeric_verbname_strings exists with a true value, then func-

tions on verbs will also accept a numeric string (e.g., "4") as a verb descriptor. The decimal integer in the string works more-or-less like the positive integers described above, but with two significant differences:

1. The numeric string is a zero-based index into verbs(); that is, in the string case, you would use the number one less than what you would use in the positive integer case.

2. When there exists a verb whose actual name looks like a decimal integer, this numeric-string notation is ambiguous; the server will in all cases assume that the reference is to the first verb in the list for which the given string could be a name, either in the normal sense or as a numeric index.

Clearly, this older mechanism is more difficult and risky to use; new code should only be written to use the current mechanism, and old code using numeric strings should be modified not to do so.

Function list verb info (obj object, str verb-desc) Function none set verb info (obj object, str verb-desc, list info)

These two functions get and set (respectively) the owner, permission bits, and name(s) for the verb as specified by verb-desc on the given object. If object is not valid, then E_INVARG is raised. If object does not define a verb as specified by verb-desc, then E_VERBNF is raised. If the programmer does not have read (write) permission on the verb in question, then verb_info() (set_verb_info()) raises E_PERM. Verb info has the following form:

{owner, perms, names} where owner is an object, perms is a string containing only characters from the set ‘r’, ‘w’, ‘x’, and ‘d’, and names is a string. This is the kind of value returned by verb_

info() and expected as the third argument to set_verb_info(). set_verb_info()

raises E_INVARG if owner is not valid, if perms contains any illegal characters, or if names is the empty string or consists entirely of spaces; it raises E_PERM if owner is not the programmer and the programmer is not a wizard.

Function list verb args (obj object, str verb-desc) Function none set verb args (obj object, str verb-desc, list args)

These two functions get and set (respectively) the direct-object, preposition, and indirect-object specifications for the verb as specified by verb-desc on the given object. If object is not valid, then E_INVARG is raised. If object does not define a verb as specified by verb-desc, then E_VERBNF is raised. If the programmer does not have read (write) permission on the verb in question, then verb_args() (set_verb_args()) raises E_PERM. Verb args specifications have the following form:

{dobj, prep, iobj} where dobj and iobj are strings drawn from the set "this", "none", and "any", and prep is a string that is either "none", "any", or one of the prepositional phrases listed much earlier in the description of verbs in the first chapter. This is the kind of value returned by verb_args() and expected as the third argument to set_verb_args(). Note that for set_verb_args(), prep must be only one of the prepositional phrases, not (as is shown in that table) a set of such phrases separated by ‘/’ characters. set_ verb_args raises E_INVARG if any of the dobj, prep, or iobj strings is illegal.

     verb_args($container, "take")
                         ⇒ {"any", "out of/from inside/from", "this"}
     set_verb_args($container, "take", {"any", "from", "this"})

Function none add verb (obj object, list info, list args)

Defines a new verb on the given object. The new verb’s owner, permission bits and name(s) are given by info in the same format as is returned by verb_info(), described above. The new verb’s direct-object, preposition, and indirect-object specifications are given by args in the same format as is returned by verb_args, described above. The new verb initially has the empty program associated with it; this program does nothing but return an unspecified value. If object is not valid, or info does not specify a valid owner and well-formed permission bits and verb names, or args is not a legitimate syntax specification, then E_INVARG is raised. If the programmer does not have write permission on object or if the owner specified by info is not the programmer and the programmer is not a wizard, then E_PERM is raised.

Function none delete verb (obj object, str verb-desc)

Removes the verb as specified by verb-desc from the given object. If object is not valid, then E_INVARG is raised. If the programmer does not have write permission on object, then E_PERM is raised. If object does not define a verb as specified by verb-desc, then E_VERBNF is raised.

Function list verb code (obj object, str verb-desc [, fully-paren [, indent]]) Function list set verb code (obj object, str verb-desc, list code)

These functions get and set (respectively) the MOO-code program associated with the verb as specified by verb-desc on object. The program is represented as a list of strings, one for each line of the program; this is the kind of value returned by verb_ code() and expected as the third argument to set_verb_code(). For verb_code(), the expressions in the returned code are usually written with the minimum-necessary parenthesization; if full-paren is true, then all expressions are fully parenthesized. Also for verb_code(), the lines in the returned code are usually not indented at all; if indent is true, each line is indented to better show the nesting of statements. If object is not valid, then E_INVARG is raised. If object does not define a verb as specified by verb-desc, then E_VERBNF is raised. If the programmer does not have read (write) permission on the verb in question, then verb_code() (set_verb_code()) raises E_PERM. If the programmer is not, in fact. a programmer, then E_PERM is raised. For set_verb_code(), the result is a list of strings, the error messages generated by the MOO-code compiler during processing of code. If the list is non-empty, then set_ verb_code() did not install code; the program associated with the verb in question is unchanged.

Function list disassemble (obj object, str verb-desc)

Returns a (longish) list of strings giving a listing of the server’s internal “compiled” form of the verb as specified by verb-desc on object. This format is not documented and may indeed change from release to release, but some programmers may nonetheless find the output of disassemble() interesting to peruse as a way to gain a deeper appreciation of how the server works. If object is not valid, then E_INVARG is raised. If object does not define a verb as specified by verb-desc, then E_VERBNF is raised. If the programmer does not have read permission on the verb in question, then disassemble() raises E_PERM.

4.4.3.5 Operations on Player Objects

Function list players ()

Returns a list of the object numbers of all player objects in the database.

Function int is player (obj object)

Returns a true value if the given object is a player object and a false value otherwise. If object is not valid, E_INVARG is raised.

Function none set player flag (obj object, value)

Confers or removes the “player object” status of the given object, depending upon the truth value of value. If object is not valid, E_INVARG is raised. If the programmer is not a wizard, then E_PERM is raised. If value is true, then object gains (or keeps) “player object” status: it will be an element of the list returned by players(), the expression is_player(object) will return true, and the server will treat a call to $do_login_command() that returns object as logging in the current connection. If value is false, the object loses (or continues to lack) “player object” status: it will not be an element of the list returned by players(), the expression is_player(object) will return false, and users cannot connect to object by name when they log into the server. In addition, if a user is connected to object at the time that it loses “player object” status, then that connection is immediately broken, just as if boot_player(object) had been called (see the description of boot_player() below).

4.4.4 Operations on Network Connections

Function list connected players ([include-all])

Returns a list of the object numbers of those player objects with currently-active connections. If include-all is provided and true, then the list includes the object numbers associated with all current connections, including ones that are outbound and/or not yet logged-in.

Function int connected seconds (obj player) Function int idle seconds (obj player)

These functions return the number of seconds that the currently-active connection to player has existed and been idle, respectively. If player is not the object number of a player object with a currently-active connection, then E_INVARG is raised.

Function none notify (obj conn, str string [, no-flush])

Enqueues string for output (on a line by itself) on the connection conn. If the programmer is not conn or a wizard, then E_PERM is raised. If conn is not a currently-active connection, then this function does nothing. Output is normally written to connections only between tasks, not during execution. The server will not queue an arbitrary amount of output for a connection; the MAX_ QUEUED_OUTPUT compilation option (in ‘options.h’) controls the limit. When an attempt is made to enqueue output that would take the server over its limit, it first tries to write as much output as possible to the connection without having to wait for the other end. If that doesn’t result in the new output being able to fit in the queue, the server starts throwing away the oldest lines in the queue until the new ouput will fit. The server remembers how many lines of output it has ‘flushed’ in this way and, when next it can succeed in writing anything to the connection, it first writes a line like

>> Network buffer overflow: X lines of output to you have been lost << where

X is the number of flushed lines. If no-flush is provided and true, then notify() never flushes any output from the queue; instead it immediately returns false. Notify() otherwise always returns true.

Function int buffered output length ([obj conn])

Returns the number of bytes currently buffered for output to the connection conn. If conn is not provided, returns the maximum number of bytes that will be buffered up for output on any connection.

Function str read ([obj conn [, non-blocking]])

Reads and returns a line of input from the connection conn (or, if not provided, from the player that typed the command that initiated the current task). If non-blocking is false or not provided, this function suspends the current task, resuming it when there is input available to be read. If non-blocking is provided and true, this function never suspends the calling task; if there is no input currently available for input, read() simply returns 0 immediately. If player is provided, then the programmer must either be a wizard or the owner of player; if player is not provided, then read() may only be called by a wizard and only in the task that was last spawned by a command from the connection in question. Otherwise, E_PERM is raised. If the given player is not currently connected and has no pending lines of input, or if the connection is closed while a task is waiting for input but before any lines of input are received, then read() raises E_INVARG. The restriction on the use of read() without any arguments preserves the following simple invariant: if input is being read from a player, it is for the task started by the last command that player typed. This invariant adds responsibility to the programmer, however. If your program calls another verb before doing a read(), then either that verb must not suspend or else you must arrange that no commands will be read from the connection in the meantime. The most straightforward way to do this is to call

     set_connection_option(player, "hold-input", 1)

before any task suspension could happen, then make all of your calls to read() and other code that might suspend, and finally call

     set_connection_option(player, "hold-input", 0)

to allow commands once again to be read and interpreted normally.

Function none force input (obj conn, str line [, at-front])

Inserts the string line as an input task in the queue for the connection conn, just as if it had arrived as input over the network. If at front is provided and true, then the new line of input is put at the front of conn’s queue, so that it will be the very next line of input processed even if there is already some other input in that queue. Raises E_INVARG if conn does not specify a current connection and E_PERM if the programmer is neither conn nor a wizard.

Function none flush input (obj conn [show-messages])

Performs the same actions as if the connection conn’s defined flush command had been received on that connection, i.e., removes all pending lines of input from conn’s queue and, if show-messages is provided and true, prints a message to conn listing the flushed lines, if any. See the chapter on server assumptions about the database for more information about a connection’s defined flush command.

Function list output delimiters (obj player)

Returns a list of two strings, the current output prefix and output suffix for player. If player does not have an active network connection, then E_INVARG is raised. If either string is currently undefined, the value "" is used instead. See the discussion of the PREFIX and SUFFIX commands in the next chapter for more information about the output prefix and suffix.

Function none boot player (obj player)

Marks for disconnection any currently-active connection to the given player. The connection will not actually be closed until the currently-running task returns or suspends, but all MOO functions (such as notify(), connected_players(), and the like) immediately behave as if the connection no longer exists. If the programmer is not either a wizard or the same as player, then E_PERM is raised. If there is no currently-active connection to player, then this function does nothing. If there was a currently-active connection, then the following verb call is made when the connection is actually closed:

     $user_disconnected(player)

It is not an error if this verb does not exist; the call is simply skipped.

Function str connection name (obj player)

Returns a network-specific string identifying the connection being used by the given player. If the programmer is not a wizard and not player, then E_PERM is raised. If player is not currently connected, then E_INVARG is raised. For the TCP/IP networking configurations, for in-bound connections, the string has the form

     "port lport from host, port port"

where lport is the decimal TCP listening port on which the connection arrived, host is either the name or decimal TCP address of the host from which the player is connected, and port is the decimal TCP port of the connection on that host. For outbound TCP/IP connections, the string has the form

     "port lport to host, port port"

where lport is the decimal local TCP port number from which the connection originated, host is either the name or decimal TCP address of the host to which the connection was opened, and port is the decimal TCP port of the connection on that host. For the System V ‘local’ networking configuration, the string is the UNIX login name of the connecting user or, if no such name can be found, something of the form

     "User #number"

where number is a UNIX numeric user ID. For the other networking configurations, the string is the same for all connections and, thus, useless.

Function none set connection option (obj conn, str option, value)

Controls a number of optional behaviors associated the connection conn. Raises E_ INVARG if conn does not specify a current connection and E_PERM if the programmer is neither conn nor a wizard. The following values for option are currently supported:

"hold-input"

If value is true, then input received on conn will never be treated as a command; instead, it will remain in the queue until retrieved by a call to

read().
"client-echo"

Send the Telnet Protocol ‘WONT ECHO’ or ‘WILL ECHO’ command, depending on whether value is true or false, respectively. For clients that support the Telnet Protocol, this should toggle whether or not the client echoes locally the characters typed by the user. Note that the server itself never echoes input characters under any circumstances. (This option is only available under the TCP/IP networking configurations.)

"binary" If value is true, then both input from and output to conn can contain arbitrary bytes. Input from a connection in binary mode is not broken into lines at all; it is delivered to either the read() function or the builtin command parser as binary strings, in whatever size chunks come back from the operating system. (See the early section on MOO value types for a description of the binary string representation.) For output to a connection in binary mode, the second argument to ‘notify()’ must be a binary string; if it is malformed, E INVARG is raised.

"flush-command"

If value is a non-empty string, then it becomes the new flush command for this connection, by which the player can flush all queued input that has not yet been processed by the server. If value is not a non-empty string, then conn is set to have no flush command at all. The default value of this option

can be set via the property $server_options.default_flush_command;

see the chapter on server assumptions about the database for details.

Function list connection options (obj conn)

Returns a list of {name, value} pairs describing the current settings of all of the allowed options for the connection conn. Raises E_INVARG if conn does not specify a current connection and E_PERM if the programmer is neither conn nor a wizard.

Function value connection option (obj conn, str name)

Returns the current setting of the option name for the connection conn. Raises E_ INVARG if conn does not specify a current connection and E_PERM if the programmer is neither conn nor a wizard.

Function obj open network connection (value, …)

Establishes a network connection to the place specified by the arguments and moreor-less pretends that a new, normal player connection has been established from there. The new connection, as usual, will not be logged in initially and will have a negative object number associated with it for use with read(), notify(), and boot_player(). This object number is the value returned by this function.

If the programmer is not a wizard or if the OUTBOUND_NETWORK compilation option was not used in building the server, then E_PERM is raised. If the network connection cannot be made for some reason, then other errors will be returned, depending upon the particular network implementation in use. For the TCP/IP network implementations (the only ones as of this writing that support outbound connections), there must be two arguments, a string naming a host (possibly using the numeric Internet syntax) and an integer specifying a TCP port. If a connection cannot be made because the host does not exist, the port does not exist, the host is not reachable or refused the connection, E_INVARG is raised. If the connection cannot be made for other reasons, including resource limitations, then E_QUOTA is raised. The outbound connection process involves certain steps that can take quite a long time, during which the server is not doing anything else, including responding to user commands and executing MOO tasks. See the chapter on server assumptions about the database for details about how the server limits the amount of time it will wait for these steps to successfully complete. It is worth mentioning one tricky point concerning the use of this function. Since the server treats the new connection pretty much like any normal player connection, it will naturally try to parse any input from that connection as commands in the usual way. To prevent this treatment, you should use set_connection_option() to set the "hold-input" option true on the connection.

Function value listen (obj object, point [, print-messages])

Create a new point at which the server will listen for network connections, just as it does normally. Object is the object whose verbs do_login_command, do_command, do_

out_of_band_command, user_connected, user_created, user_reconnected, user_
disconnected, and user_client_disconnected will be called at appropriate points,

just as these verbs are called on #0 for normal connections. (See the chapter on server assumptions about the database for the complete story on when these functions are called.) Point is a network-configuration-specific parameter describing the listening point. If print-messages is provided and true, then the various database-configurable messages (also detailed in the chapter on server assumptions) will be printed on connections received at the new listening point. Listen() returns canon, a ‘canonicalized’ version of point, with any configuration-specific defaulting or aliasing accounted for. This raises E_PERM if the programmer is not a wizard, E_INVARG if object is invalid or there is already a listening point described by point, and E_QUOTA if some networkconfiguration-specific error occurred. For the TCP/IP configurations, point is a TCP port number on which to listen and canon is equal to point unless point is zero, in which case canon is a port number assigned by the operating system. For the local multi-user configurations, point is the UNIX file name to be used as the connection point and canon is always equal to point. In the single-user configuration, the can be only one listening point at a time; point can be any value at all and canon is always zero.

Function none unlisten (canon)

Stop listening for connections on the point described by canon, which should be the second element of some element of the list returned by listeners(). Raises E_PERM if the programmer is not a wizard and E_INVARG if there does not exist a listener with that description.

Function list listeners ()

Returns a list describing all existing listening points, including the default one set up automatically by the server when it was started (unless that one has since been destroyed by a call to unlisten()). Each element of the list has the following form:

{object, canon, print-messages} where object is the first argument given in the call to listen() to create this listening point, print-messages is true if the third argument in that call was provided and true, and canon was the value returned by that call. (For the initial listening point, object is #0, canon is determined by the command-line arguments or a network-configuration-specific default, and print-messages is true.)

Please note that there is nothing special about the initial listening point created by the server when it starts; you can use unlisten() on it just as if it had been created by listen(). This can be useful; for example, under one of the TCP/IP configurations, you might start up your server on some obscure port, say 12345, connect to it by yourself for a while, and then open it up to normal users by evaluating the statments

unlisten(12345); listen(#0, 7777, 1)

4.4.5 Operations Involving Times and Dates

Function int time ()

Returns the current time, represented as the number of seconds that have elapsed since midnight on 1 January 1970, Greenwich Mean Time.

Function str ctime ([int time])

Interprets time as a time, using the same representation as given in the description of time(), above, and converts it into a 28-character, human-readable string in the following format:

     Mon Aug 13 19:13:20 1990 PDT

If the current day of the month is less than 10, then an extra blank appears between the month and the day:

     Mon Apr  1 14:10:43 1991 PST

If time is not provided, then the current time is used. Note that ctime() interprets time for the local time zone of the computer on which the MOO server is running.

4.4.6 MOO-Code Evaluation and Task Manipulation

Function none raise (code [, str message [, value]])

Raises code as an error in the same way as other MOO expressions, statements, and functions do. Message, which defaults to the value of tostr(code), and value, which defaults to zero, are made available to any try-except statements that catch the error. If the error is not caught, then message will appear on the first line of the traceback printed to the user.

Function value call function (str func-name, arg, …)

Calls the built-in function named func-name, passing the given arguments, and returns whatever that function returns. Raises E_INVARG if func-name is not recognized as the name of a known built-in function. This allows you to compute the name of the function to call and, in particular, allows you to write a call to a built-in function that may or may not exist in the particular version of the server you’re using.

Function list function info ([str name])

Returns descriptions of the built-in functions available on the server. If name is provided, only the description of the function with that name is returned. If name is omitted, a list of descriptions is returned, one for each function available on the server. Raised E_INVARG if name is provided but no function with that name is available on the server.

Each function description is a list of the following form:

{name, min-args, max-args, types

where name is the name of the built-in function, min-args is the minimum number of arguments that must be provided to the function, max-args is the maximum number of arguments that can be provided to the function or -1 if there is no maximum, and types is a list of max-args integers (or min-args if max-args is -1), each of which represents the type of argument required in the corresponding position. Each type number is as would be returned from the typeof() built-in function except that -1 indicates that any type of value is acceptable and -2 indicates that either integers or floating-point numbers may be given. For example, here are several entries from the list:

     {"listdelete", 2, 2, {4, 0}}
     {"suspend", 0, 1, {0}}
     {"server_log", 1, 2, {2, -1}}
     {"max", 1, -1, {-2}}
     {"tostr", 0, -1, {}}

Listdelete() takes exactly 2 arguments, of which the first must be a list (LIST == 4) and the second must be an integer (INT == 0). Suspend() has one optional argument that, if provided, must be an integer. Server_log() has one required argument that must be a string (STR == 2) and one optional argument that, if provided, may be of any type. Max() requires at least one argument but can take any number above that, and the first argument must be either an integer or a floating-point number; the type(s) required for any other arguments can’t be determined from this description. Finally, tostr() takes any number of arguments at all, but it can’t be determined from this description which argument types would be acceptable in which positions.

Function list eval (str string)

The MOO-code compiler processes string as if it were to be the program associated with some verb and, if no errors are found, that fictional verb is invoked. If the programmer is not, in fact, a programmer, then E_PERM is raised. The normal result of calling eval() is a two element list. The first element is true if there were no compilation errors and false otherwise. The second element is either the result returned from the fictional verb (if there were no compilation errors) or a list of the compiler’s error messages (otherwise).

When the fictional verb is invoked, the various built-in variables have values as shown below:

player the same as in the calling verb

     this      #-1

caller the same as the initial value of this in the calling verb

     args      {}
     argstr    ""
     verb      ""
     dobjstr   ""
     dobj      #-1
     prepstr   ""
     iobjstr   ""
     iobj      #-1

The fictional verb runs with the permissions of the programmer and as if its ‘d’ permissions bit were on.

     eval("return 3 + 4;") ⇒ {1, 7}

Function none set task perms (obj who)

Changes the permissions with which the currently-executing verb is running to be those of who. If the programmer is neither who nor a wizard, then E_PERM is raised.

Note: This does not change the owner of the currently-running verb, only the permissions of this particular invocation. It is used in verbs owned by wizards to make themselves run with lesser (usually non-wizard) permissions.

Function obj caller perms ()

Returns the permissions in use by the verb that called the currently-executing verb. If the currently-executing verb was not called by another verb (i.e., it is the first verb called in a command or server task), then caller_perms() returns #-1.

Function int ticks left () Function int seconds left ()

These two functions return the number of ticks or seconds (respectively) left to the current task before it will be forcibly terminated. These are useful, for example, in deciding when to call ‘suspend()’ to continue a long-lived computation.

Function int task id ()

Returns the non-zero, non-negative integer identifier for the currently-executing task. Such integers are randomly selected for each task and can therefore safely be used in circumstances where unpredictability is required.

Function value suspend ([int seconds])

Suspends the current task, and resumes it after at least seconds seconds. (If seconds is not provided, the task is suspended indefinitely; such a task can only be resumed by use of the resume() function.) When the task is resumed, it will have a full quota of ticks and seconds. This function is useful for programs that run for a long time or require a lot of ticks. If seconds is negative, then E_INVARG is raised. Suspend() returns zero unless it was resumed via resume(), in which case it returns the second argument given to that function. In some sense, this function forks the ‘rest’ of the executing task. However, there is a major difference between the use of ‘suspend(seconds)’ and the use of the ‘fork (seconds)’. The ‘fork’ statement creates a new task (a forked task) while the currentlyrunning task still goes on to completion, but a suspend() suspends the currentlyrunning task (thus making it into a suspended task). This difference may be best explained by the following examples, in which one verb calls another:

     .program   #0:caller_A
     #0.prop = 1;
     #0:callee_A();
     #0.prop = 2;
     .
     .program   #0:callee_A
     fork(5)
       #0.prop = 3;
     endfork
     .
     .program   #0:caller_B
     #0.prop = 1;
     #0:callee_B();
     #0.prop = 2;
     .
     .program   #0:callee_B
     suspend(5);
     #0.prop = 3;
     .

Consider #0:caller_A, which calls #0:callee_A. Such a task would assign 1 to #0.prop, call #0:callee_A, fork a new task, return to #0:caller_A, and assign 2 to #0.prop, ending this task. Five seconds later, if the forked task had not been killed, then it would begin to run; it would assign 3 to #0.prop and then stop. So, the final value of #0.prop (i.e., the value after more than 5 seconds) would be 3. Now consider #0:caller_B, which calls #0:callee_B instead of #0:callee_A. This task would assign 1 to #0.prop, call #0:callee_B, and suspend. Five seconds later, if the suspended task had not been killed, then it would resume; it would assign 3 to #0.prop, return to #0:caller_B, and assign 2 to #0.prop, ending the task. So, the final value of #0.prop (i.e., the value after more than 5 seconds) would be 2. A suspended task, like a forked task, can be described by the queued_tasks() function and killed by the kill_task() function. Suspending a task does not change its task id. A task can be suspended again and again by successive calls to suspend(). By default, there is no limit to the number of tasks any player may suspend, but such a limit can be imposed from within the database. See the chapter on server assumptions about the database for details.

Function none resume (int task-id [, value])

Immediately ends the suspension of the suspended task with the given task-id; that task’s call to suspend() will return value, which defaults to zero. If value is of type ERR, it will be raised, rather than returned, in the suspended task. Resume() raises E_INVARG if task-id does not specify an existing suspended task and E_PERM if the programmer is neither a wizard nor the owner of the specified task.

Function list queue info ([obj player])

If player is omitted, returns a list of object numbers naming all players that currently have active task queues inside the server. If player is provided, returns the number of background tasks currently queued for that user. It is guaranteed that queue_info(X) will return zero for any X not in the result of queue_info().

Function list queued tasks ()

Returns information on each of the background tasks (i.e., forked, suspended or reading) owned by the programmer (or, if the programmer is a wizard, all queued tasks). The returned value is a list of lists, each of which encodes certain information about a particular queued task in the following format:

{task-id, start-time, x, y,

programmer, verb-loc, verb-name, line, this} where task-id is an integer identifier for this queued task, start-time is the time after which this task will begin execution (in time() format), x and y are obsolete values that are no longer interesting, programmer is the permissions with which this task will begin execution (and also the player who owns this task), verb-loc is the object on which the verb that forked this task was defined at the time, verb-name is that name of that verb, line is the number of the first line of the code in that verb that this task will execute, and this is the value of the variable ‘this’ in that verb. For reading tasks, start-time is -1. The x and y fields are now obsolete and are retained only for backward-compatibility reasons. They may be reused for new purposes in some future version of the server.

Function none kill task (int task-id)

Removes the task with the given task-id from the queue of waiting tasks. If the programmer is not the owner of that task and not a wizard, then E_PERM is raised. If there is no task on the queue with the given task-id, then E_INVARG is raised.

Function list callers ([include-line-numbers])

Returns information on each of the verbs and built-in functions currently waiting to resume execution in the current task. When one verb or function calls another verb or function, execution of the caller is temporarily suspended, pending the called verb or function returning a value. At any given time, there could be several such pending verbs and functions: the one that called the currently executing verb, the verb or function that called that one, and so on. The result of callers() is a list, each element of which gives information about one pending verb or function in the following format:

{this, verb-name, programmer, verb-loc, player, line-number} For verbs, this is the initial value of the variable ‘this’ in that verb, verb-name is the name used to invoke that verb, programmer is the player with whose permissions that verb is running, verb-loc is the object on which that verb is defined, player is the initial value of the variable ‘player’ in that verb, and line-number indicates which line of the verb’s code is executing. The line-number element is included only if the include-line-numbers argument was provided and true. For functions, this, programmer, and verb-loc are all #-1, verb-name is the name of the function, and line-number is an index used internally to determine the current state of the built-in function. The simplest correct test for a built-in function entry is

     (VERB-LOC == #-1  &&  PROGRAMMER == #-1  &&  VERB-NAME != "")

The first element of the list returned by callers() gives information on the verb that called the currently-executing verb, the second element describes the verb that called that one, and so on. The last element of the list describes the first verb called in this task.

Function list task stack (int task-id [, include-line-numbers])

Returns information like that returned by the callers() function, but for the suspended task with the given task-id; the include-line-numbers argument has the same meaning as in callers(). Raises E_INVARG if task-id does not specify an existing suspended task and E_PERM if the programmer is neither a wizard nor the owner of the specified task.

4.4.7 Administrative Operations

Function str server version ()

Returns a string giving the version number of the running MOO server.

Function none server log (str message [, is-error])

The text in message is sent to the server log with a distinctive prefix (so that it can be distinguished from server-generated messages). If the programmer is not a wizard, then E_PERM is raised. If is-error is provided and true, then message is marked in the server log as an error.

Function obj renumber (obj object)

The object number of the object currently numbered object is changed to be the least nonnegative object number not currently in use and the new object number is returned. If object is not valid, then E_INVARG is raised. If the programmer is not a wizard, then E_PERM is raised. If there are no unused nonnegative object numbers less than object, then object is returned and no changes take place. The references to object in the parent/children and location/contents hierarchies are updated to use the new object number, and any verbs, properties and/or objects owned by object are also changed to be owned by the new object number. The latter operation can be quite time consuming if the database is large. No other changes to the database are performed; in particular, no object references in property values or verb code are updated. This operation is intended for use in making new versions of the LambdaCore database from the then-current LambdaMOO database, and other similar situations. Its use requires great care.

Function none reset max object ()

The server’s idea of the highest object number ever used is changed to be the highest object number of a currently-existing object, thus allowing reuse of any higher numbers that refer to now-recycled objects. If the programmer is not a wizard, then E_PERM is raised. This operation is intended for use in making new versions of the LambdaCore database from the then-current LambdaMOO database, and other similar situations. Its use requires great care.

Function list memory usage ()

On some versions of the server, this returns statistics concerning the server consumption of system memory. The result is a list of lists, each in the following format:

{block-size, nused, nfree} where block-size is the size in bytes of a particular class of memory fragments, nused is the number of such fragments currently in use in the server, and nfree is the number of such fragments that have been reserved for use but are currently free. On servers for which such statistics are not available, memory_usage() returns {}. The compilation option USE_GNU_MALLOC controls whether or not statistics are available; if the option is not provided, statistics are not available.

Function none dump database ()

Requests that the server checkpoint the database at its next opportunity. It is not normally necessary to call this function; the server automatically checkpoints the database at regular intervals; see the chapter on server assumptions about the database for details. If the programmer is not a wizard, then E_PERM is raised.

Function int db disk size ()

Returns the total size, in bytes, of the most recent full representation of the database as one or more disk files. Raises E_QUOTA if, for some reason, no such on-disk representation is currently available.

Function none shutdown ([str message])

Requests that the server shut itself down at its next opportunity. Before doing so, a notice (incorporating message, if provided) is printed to all connected players. If the programmer is not a wizard, then E_PERM is raised.

5 Server Commands and Database Assumptions

This chapter describes all of the commands that are built into the server and every property and verb in the database specifically accessed by the server. Aside from what is listed here, no assumptions are made by the server concerning the contents of the database.

5.1 Built-in Commands

As was mentioned in the chapter on command parsing, there are five commands whose inter-

pretation is fixed by the server: PREFIX, OUTPUTPREFIX, SUFFIX, OUTPUTSUFFIX, and .program.

The first four of these are intended for use by programs that connect to the MOO, so-called ‘client’ programs. The .program command is used by programmers to associate a MOO program with a particular verb. The server can, in addition, recognize a sixth special command on any or all connections, the flush command.

The server also performs special processing on command lines that begin with certain punctuation characters.

This section discusses these built-in pieces of the command-interpretation process.

5.1.1 Command-Output Delimiters

Every MOO network connection has associated with it two strings, the output prefix and the output suffix. Just before executing a command typed on that connection, the server prints the output prefix, if any, to the player. Similarly, just after finishing the command, the output suffix, if any, is printed to the player. Initially, these strings are not defined, so no extra printing takes place.

The PREFIX and SUFFIX commands are used to set and clear these strings. They have the following simple syntax:

PREFIX output-prefix SUFFIX output-suffix

That is, all text after the command name and any following spaces is used as the new value of the appropriate string. If there is no non-blank text after the command string, then the corresponding string is cleared. For compatibility with some general MUD client programs, the server also recognizes OUTPUTPREFIX as a synonym for PREFIX and OUTPUTSUFFIX as a synonym for SUFFIX.

These commands are intended for use by programs connected to the MOO, so that they can issue MOO commands and reliably determine the beginning and end of the resulting output. For example, one editor-based client program sends this sequence of commands on occasion:

     PREFIX >>MOO-Prefix<<
     SUFFIX >>MOO-Suffix<<
     @list object:verb without numbers
     PREFIX
     SUFFIX

The effect of which, in a LambdaCore-derived database, is to print out the code for the named verb preceded by a line containing only ‘>>MOO-Prefix<<’ and followed by a line containing only ‘>>MOO-Suffix<<’. This enables the editor to reliably extract the program text from the MOO output and show it to the user in a separate editor window. There are many other possible uses.

The built-in function output_delimiters() can be used by MOO code to find out the output prefix and suffix currently in effect on a particular network connection.

5.1.2 Programming

The .program command is a common way for programmers to associate a particular MOO-code program with a particular verb. It has the following syntax:

.program object:verb

…several lines of MOO code…

     .

That is, after typing the .program command, then all lines of input from the player are considered to be a part of the MOO program being defined. This ends as soon as the player types a line containing only a dot (‘.’). When that line is received, the accumulated MOO program is checked for proper MOO syntax and, if correct, associated with the named verb.

If, at the time the line containing only a dot is processed, (a) the player is not a programmer,

(b) the player does not have write permission on the named verb, or (c) the property $server_ options.protect_set_verb_code exists and has a true value and the player is not a wizard, then an error message is printed and the named verb’s program is not changed.

In the .program command, object may have one of three forms:

• The name of some object visible to the player. This is exactly like the kind of matching done by the server for the direct and indirect objects of ordinary commands. See the chapter on command parsing for details. Note that the special names ‘me’ and ‘here’ may be used.

• An object number, in the form #number.

• A system property (that is, a property on #0), in the form $name. In this case, the current value of #0.name must be a valid object.

5.1.3 Flushing Unprocessed Input

It sometimes happens that a user changes their mind about having typed one or more lines of input and would like to ‘untype’ them before the server actually gets around to processing them. If they react quickly enough, they can type their connection’s defined flush command; when the server first reads that command from the network, it immediately and completely flushes any as-yet unprocessed input from that user, printing a message to the user describing just which lines of input were discarded, if any.

Fine point: The flush command is handled very early in the server’s processing of a line of input, before the line is entered into the task queue for the connection and well before it is parsed into words like other commands. For this reason, it must be typed exactly as it was defined, alone on the line, without quotation marks, and without any spaces before or after it.

When a connection is first accepted by the server, it is given an initial flush command setting taken from the current default. This initial setting can be changed later using the set_connection_

option() command.

By default, each connection is initially given ‘.flush’ as its flush command. If the property $server_options.default_flush_command exists, then its value overrides this default. If $server_options.default_flush_command is a non-empty string, then that string is the flush command for all new connections; otherwise, new connections are initially given no flush command at all.

5.1.4 Initial Punctuation in Commands

The server interprets command lines that begin with any of the following characters specially:

     "        :        ;

Before processing the command, the initial punctuation character is replaced by the corresponding word below, followed by a space:

     say      emote    eval

For example, the command line

     "Hello, there.

is transformed into

     say Hello, there.

before parsing.

5.2 Server Assumptions About the Database

There are a small number of circumstances under which the server directly and specifically accesses a particular verb or property in the database. This section gives a complete list of such circumstances.

5.2.1 Server Options Set in the Database

Many optional behaviors of the server can be controlled from within the database by creating the property #0.server_options (also known as $server_options), assigning as its value a valid object number, and then defining various properties on that object. At a number of times, the server checks for whether the property $server_options exists and has an object number as its value. If so, then the server looks for a variety of other properties on that $server_options object and, if they exist, uses their values to control how the server operates.

The specific properties searched for are each described in the appropriate section below, but here is a brief list of all of the relevant properties for ease of reference:

bg_seconds

The number of seconds allotted to background tasks.

bg_ticks The number of ticks allotted to background tasks.

connect_timeout

The maximum number of seconds to allow an un-logged-in in-bound connection to remain open.

default_flush_command

The initial setting of each new connection’s flush command.

fg_seconds

The number of seconds allotted to foreground tasks.

fg_ticks The number of ticks allotted to foreground tasks.

max_stack_depth

The maximum number of levels of nested verb calls.

name_lookup_timeout

The maximum number of seconds to wait for a network hostname/address lookup.

outbound_connect_timeout

The maximum number of seconds to wait for an outbound network connection to successfully open.

protect_property

Restrict reading of built-in property to wizards.

protect_function

Restrict use of built-in function to wizards.

support_numeric_verbname_strings

Enables use of an obsolete verb-naming mechanism.

5.2.2 Server Messages Set in the Database

There are a number of circumstances under which the server itself generates messages on network connections. Most of these can be customized or even eliminated from within the database. In each such case, a property on $server_options is checked at the time the message would be printed. If the property does not exist, a default message is printed. If the property exists and its value is not a string or a list containing strings, then no message is printed at all. Otherwise, the string(s) are printed in place of the default message, one string per line. None of these messages are ever printed on an outbound network connection created by the function open_network_connection().

The following list covers all of the customizable messages, showing for each the name of the relevant property on $server_options, the default message, and the circumstances under which the message is printed:

boot_msg = "*** Disconnected ***"

The function boot_player() was called on this connection.

connect_msg = "*** Connected ***"

The user object that just logged in on this connection existed before $do_login_ command() was called.

create_msg = "*** Created ***"

The user object that just logged in on this connection did not exist before $do_login_ command() was called.

recycle_msg = "*** Recycled ***"

The logged-in user of this connection has been recycled or renumbered (via the renumber() function).

redirect_from_msg = "*** Redirecting connection to new port ***"

The logged-in user of this connection has just logged in on some other connection.

redirect_to_msg = "*** Redirecting old connection to this port ***"

The user who just logged in on this connection was already logged in on some other connection.

server_full_msg

Default:

               *** Sorry, but the server cannot accept any more connections right now.
               *** Please try again later.

This connection arrived when the server really couldn’t accept any more connections, due to running out of a critical operating system resource.

timeout_msg = "*** Timed-out waiting for login. ***"

This in-bound network connection was idle and un-logged-in for at least CONNECT_ TIMEOUT seconds (as defined in the file ‘options.h’ when the server was compiled).

Fine point: If the network connection in question was received at a listening point (established by the ‘listen()’ function) handled by an object obj other than #0, then system messages for that connection are looked for on obj.server_options; if that property does not exist, then $server_options is used instead.

5.2.3 Checkpointing the Database

The server maintains the entire MOO database in main memory, not on disk. It is therefore necessary for it to dump the database to disk if it is to persist beyond the lifetime of any particular server execution. The server is careful to dump the database just before shutting down, of course, but it is also prudent for it to do so at regular intervals, just in case something untoward happens.

To determine how often to make these checkpoints of the database, the server consults the value of #0.dump_interval. If it exists and its value is an integer greater than or equal to 60, then it is taken as the number of seconds to wait between checkpoints; otherwise, the server makes a new checkpoint every 3600 seconds (one hour). If the value of #0.dump_interval implies that the next checkpoint should be scheduled at a time after 3:14:07 a.m. on Tuesday, January 19, 2038, then the server instead uses the default value of 3600 seconds in the future.

The decision about how long to wait between checkpoints is made again immediately after each one begins. Thus, changes to #0.dump_interval will take effect after the next checkpoint happens.

Whenever the server begins to make a checkpoint, it makes the following verb call:

     $checkpoint_started()

When the checkpointing process is complete, the server makes the following verb call:

     $checkpoint_finished(success)

where success is true if and only if the checkpoint was successfully written on the disk. Checkpointing can fail for a number of reasons, usually due to exhaustion of various operating system resources such as virtual memory or disk space. It is not an error if either of these verbs does not exist; the corresponding call is simply skipped.

5.2.4 Accepting and Initiating Network Connections

When the server first accepts a new, incoming network connection, it is given the low-level network address of computer on the other end. It immediately attempts to convert this address into the human-readable host name that will be entered in the server log and returned by the connection_name() function. This conversion can, for the TCP/IP networking configurations, involve a certain amount of communication with remote name servers, which can take quite a long time and/or fail entirely. While the server is doing this conversion, it is not doing anything else at all; in particular, it it not responding to user commands or executing MOO tasks.

By default, the server will wait no more than 5 seconds for such a name lookup to succeed; after that, it behaves as if the conversion had failed, using instead a printable representation of the low-level address. If the property name_lookup_timeout exists on $server_options and has an integer as its value, that integer is used instead as the timeout interval.

When the open_network_connection() function is used, the server must again do a conversion, this time from the host name given as an argument into the low-level address necessary for actually opening the connection. This conversion is subject to the same timeout as in the in-bound case; if the conversion does not succeed before the timeout expires, the connection attempt is aborted and

open_network_connection() raises E_QUOTA.

After a successful conversion, though, the server must still wait for the actual connection to be accepted by the remote computer. As before, this can take a long time during which the server is again doing nothing else. Also as before, the server will by default wait no more than 5 seconds for the connection attempt to succeed; if the timeout expires, open_network_connection() again raises E_QUOTA. This default timeout interval can also be overridden from within the database, by defining the property outbound_connect_timeout on $server_options with an integer as its value.

5.2.5 Associating Network Connections with Players

When a network connection is first made to the MOO, it is identified by a unique, negative object number. Such a connection is said to be un-logged-in and is not yet associated with any MOO player object.

Each line of input on an un-logged-in connection is first parsed into words in the usual way (see the chapter on command parsing for details) and then these words are passed as the arguments in a call to the verb $do_login_command(). For example, the input line

     connect Munchkin frebblebit

would result in the following call being made:

     $do_login_command("connect", "Munchkin", "frebblebit")

In that call, the variable player will have as its value the negative object number associated with the appropriate network connection. The functions notify() and boot_player() can be used with such object numbers to send output to and disconnect un-logged-in connections. Also, the variable argstr will have as its value the unparsed command line as received on the network connection.

If $do_login_command() returns a valid player object and the connection is still open, then the connection is considered to have logged into that player. The server then makes one of the following verbs calls, depending on the player object that was returned:

     $user_created(player)
     $user_connected(player)
     $user_reconnected(player)

The first of these is used if the returned object number is greater than the value returned by the max_ object() function before $do_login_command() was invoked, that is, it is called if the returned object appears to have been freshly created. If this is not the case, then one of the other two verb calls is used. The $user_connected() call is used if there was no existing active connection for the returned player object. Otherwise, the $user_reconnected() call is used instead.

Fine point: If a user reconnects and the user’s old and new connections are on two different listening points being handled by different objects (see the description of the listen() function for more details), then user_client_disconnected is called for the old connection and user_connected for the new one.

If an in-bound network connection does not successfully log in within a certain period of time, the server will automatically shut down the connection, thereby freeing up the resources associated with maintaining it. Let L be the object handling the listening point on which the connection was received (or #0 if the connection came in on the initial listening point). To discover the timeout period, the server checks on L.server_options or, if it doesn’t exist, on $server_options for a connect_timeout property. If one is found and its value is a positive integer, then that’s the number of seconds the server will use for the timeout period. If the connect_timeout property exists but its value isn’t a positive integer, then there is no timeout at all. If the property doesn’t exist, then the default timeout is 300 seconds.

When any network connection (even an un-logged-in or outbound one) is terminated, by either the server or the client, then one of the following two verb calls is made:

     $user_disconnected(player)
     $user_client_disconnected(player)

The first is used if the disconnection is due to actions taken by the server (e.g., a use of the boot_ player() function or the un-logged-in timeout described above) and the second if the disconnection was initiated by the client side.

It is not an error if any of these five verbs do not exist; the corresponding call is simply skipped. Note: Only one network connection can be controlling a given player object at a given time; should a second connection attempt to log in as that player, the first connection is unceremoniously closed (and $user_reconnected() called, as described above). This makes it easy to recover from various kinds of network problems that leave connections open but inaccessible.

When the network connection is first established, the null command is automatically entered by the server, resulting in an initial call to $do_login_command() with no arguments. This signal can be used by the verb to print out a welcome message, for example.

Warning: If there is no $do_login_command() verb defined, then lines of input from un-logged-in connections are simply discarded. Thus, it is necessary for any database to include a suitable definition for this verb.

5.2.6 Out-of-Band Commands

It is possible to compile the server with an option defining an out-of-band prefix for commands. This is a string that the server will check for at the beginning of every line of input from players, regardless of whether or not those players are logged in and regardless of whether or not reading tasks are waiting for input from those players. If a given line of input begins with the defined out-of-band prefix (leading spaces, if any, are not stripped before testing), then it is not treated as a normal command or as input to any reading task. Instead, the line is parsed into a list of words in the usual way and those words are given as the arguments in a call to $do_out_of_ band_command(). For example, if the out-of-band prefix were defined to be ‘#$#’, then the line of input

     #$# client-type fancy

would result in the following call being made in a new server task:

     $do_out_of_band_command("#$#", "client-type", "fancy")

During the call to $do_out_of_band_command(), the variable player is set to the object number representing the player associated with the connection from which the input line came. Of course, if that connection has not yet logged in, the object number will be negative. Also, the variable argstr will have as its value the unparsed input line as received on the network connection.

Out-of-band commands are intended for use by fancy client programs that may generate asynchronous events of which the server must be notified. Since the client cannot, in general, know the state of the player’s connection (logged-in or not, reading task or not), out-of-band commands provide the only reliable client-to-server communications channel.

5.2.7 The First Tasks Run By the Server

Whenever the server is booted, there are a few tasks it runs right at the beginning, before accepting connections or getting the value of #0.dump_interval to schedule the first checkpoint (see below for more information on checkpoint scheduling).

First, the server calls $user_disconnected() once for each user who was connected at the time the database file was written; this allows for any cleaning up that’s usually done when users disconnect (e.g., moving their player objects back to some ‘home’ location, etc.).

Next, it checks for the existence of the verb $server_started(). If there is such a verb, then the server runs a task invoking that verb with no arguments and with player equal to #-1. This is useful for carefully scheduling checkpoints and for re-initializing any state that is not properly represented in the database file (e.g., re-opening certain outbound network connections, clearing out certain tables, etc.).

5.2.8 Controlling the Execution of Tasks

As described earlier, in the section describing MOO tasks, the server places limits on the number of seconds for which any task may run continuously and the number of “ticks,” or low-level operations, any task may execute in one unbroken period. By default, foreground tasks may use 30,000 ticks and five seconds, and background tasks may use 15,000 ticks and three seconds. These defaults can be overridden from within the database by defining any or all of the following properties on $server_options and giving them integer values:

bg_seconds

The number of seconds allotted to background tasks.

bg_ticks The number of ticks allotted to background tasks.

fg_seconds

The number of seconds allotted to foreground tasks.

fg_ticks The number of ticks allotted to foreground tasks.

The server ignores the values of fg_ticks and bg_ticks if they are less than 100 and similarly ignores fg_seconds and bg_seconds if their values are less than 1. This may help prevent utter disaster should you accidentally give them uselessly-small values.

Recall that command tasks and server tasks are deemed foreground tasks, while forked, suspended, and reading tasks are defined as background tasks. The settings of these variables take effect only at the beginning of execution or upon resumption of execution after suspending or reading.

The server also places a limit on the number of levels of nested verb calls, raising E_MAXREC from a verb-call expression if the limit is exceeded. The limit is 50 levels by default, but this can be increased from within the database by defining the max_stack_depth property on $server_ options and giving it an integer value greater than 50. The maximum stack depth for any task is set at the time that task is created and cannot be changed thereafter. This implies that suspended tasks, even after being saved in and restored from the DB, are not affected by later changes to $server options.max stack depth.

Finally, the server can place a limit on the number of forked or suspended tasks any player can have queued at a given time. Each time a fork statement or a call to suspend() is executed in some verb, the server checks for a property named queued_task_limit on the programmer. If that property exists and its value is a non-negative integer, then that integer is the limit. Otherwise, if $server_options.queued_task_limit exists and its value is a non-negative integer, then that’s the limit. Otherwise, there is no limit. If the programmer already has a number of queued tasks that is greater than or equal to the limit, E_QUOTA is raised instead of either forking or suspending. Reading tasks are affected by the queued-task limit.

5.2.9 Controlling the Handling of Aborted Tasks

The server will abort the execution of tasks for either of two reasons:

1. an error was raised within the task but not caught, or

2. the task exceeded the limits on ticks and/or seconds.

In each case, after aborting the task, the server attempts to call a particular handler verb within the database to allow code there to handle this mishap in some appropriate way. If this verb call suspends or returns a true value, then it is considered to have handled the situation completely and no further processing will be done by the server. On the other hand, if the handler verb does not exist, or if the call either returns a false value without suspending or itself is aborted, the server takes matters into its own hands.

First, an error message and a MOO verb-call stack traceback are printed to the player who typed the command that created the original aborted task, explaining why the task was aborted and where in the task the problem occurred. Then, if the call to the handler verb was itself aborted, a second error message and traceback are printed, describing that problem as well. Note that if the handler-verb call itself is aborted, no further ‘nested’ handler calls are made; this policy prevents what might otherwise be quite a vicious little cycle.

The specific handler verb, and the set of arguments it is passed, differs for the two causes of aborted tasks.

If an error is raised and not caught, then the verb-call

$handle_uncaught_error(code, msg, value, traceback, formatted) is made, where code, msg, value, and traceback are the values that would have been passed to a handler in a try-except statement and formatted is a list of strings being the lines of error and traceback output that will be printed to the player if $handle_uncaught_error returns false without suspending.

If a task runs out of ticks or seconds, then the verb-call

$handle_task_timeout(resource, traceback, formatted)

is made, where resource is the appropriate one of the strings "ticks" or "seconds", and traceback and formatted are as above.

5.2.10 Matching in Command Parsing

In the process of matching the direct and indirect object strings in a command to actual objects, the server uses the value of the aliases property, if any, on each object in the contents of the player and the player’s location. For complete details, see the chapter on command parsing.

5.2.11 Restricting Access to Built-in Properties and Functions

Whenever verb code attempts to read the value of a built-in property prop on any object, the server checks to see if the property $server_options.protect_prop exists and has a true value. If so, then E_PERM is raised if the programmer is not a wizard.

Whenever verb code calls a built-in function func() and the caller is not the object #0, the server checks to see if the property $server_options.protect_func exists and has a true value. If so, then the server next checks to see if the verb $bf_func() exists; if that verb exists, then the server calls it instead of the built-in function, returning or raising whatever that verb returns or raises. If the $bf_func() does not exist and the programmer is not a wizard, then the server immediately raises E_PERM, without actually calling the function. Otherwise (if the caller is #0, if $server_options.protect_func either doesn’t exist or has a false value, or if $bf_func() exists but the programmer is a wizard), then the built-in function is called normally.

5.2.12 Creating and Recycling Objects

Whenever the create() function is used to create a new object, that object’s initialize verb, if any, is called with no arguments. The call is simply skipped if no such verb is defined on the object.

Symmetrically, just before the recycle() function actually destroys an object, the object’s recycle verb, if any, is called with no arguments. Again, the call is simply skipped if no such verb is defined on the object.

Both create() and recycle() check for the existence of an ownership_quota property on the owner of the newly-created or -destroyed object. If such a property exists and its value is an integer, then it is treated as a quota on object ownership. Otherwise, the following two paragraphs do not apply.

The create() function checks whether or not the quota is positive; if so, it is reduced by one and stored back into the ownership_quota property on the owner. If the quota is zero or negative, the quota is considered to be exhausted and create() raises E_QUOTA.

The recycle() function increases the quota by one and stores it back into the ownership_quota property on the owner.

5.2.13 Object Movement

During evaluation of a call to the move() function, the server can make calls on the accept and enterfunc verbs defined on the destination of the move and on the exitfunc verb defined on the source. The rules and circumstances are somewhat complicated and are given in detail in the description of the move() function.

5.2.14 Temporarily Enabling Obsolete Server Features

   If the property $server_options.support_numeric_verbname_strings exists and has a true

value, then the server supports a obsolete mechanism for less ambiguously referring to specific verbs in various built-in functions. For more details, see the discussion given just following the description of the verbs() function.

Chapter 6: Function Index 78

6 Function Index

A

abs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 acos. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 add property . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 add verb . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 asin. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 atan. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40

B

binary hash . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 boot player . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57 buffered output length. . . . . . . . . . . . . . . . . . . . . . . . . . . 56

C

call function. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 caller perms . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62 callers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64 ceil. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 children . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 chparent . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 clear property . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52 connected players . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 connected seconds . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 connection name . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57 connection option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 connection options . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 cos . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 cosh. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 create . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48 crypt. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 ctime. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60

D

db disk size . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66 decode binary. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 delete property . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52 delete verb . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 disassemble. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 dump database. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66

E

encode binary. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 equal. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 eval. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 exp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40

F

floatstr(float . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 floor. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 flush input . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57 force input . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56 function info. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61

I

idle seconds . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 index. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 is clear property . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52 is member . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 is player . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55

K

kill task . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64

L

length . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41, 47 listappend . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 listdelete . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 listen . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 listeners . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 listinsert . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 listset . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 log . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 log10. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40

M

match. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 max . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 max object . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 memory usage . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66 min . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 move. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50

N

notify . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55

O

object bytes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 open network connection . . . . . . . . . . . . . . . . . . . . . . . . . 58 output delimiters . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57

Chapter 6: Function Index 79

P

parent . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 pass. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 players . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 properties . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 property info. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51

Q

queue info . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64 queued tasks . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64

R

raise. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 random . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 read. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56 recycle . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 renumber . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 reset max object. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66 resume . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64 rindex . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 rmatch . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42

S

seconds left . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62 server log . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 server version . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 set connection option . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 set player flag. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 set property info . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 set task perms . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62 set verb args . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 set verb code . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 set verb info . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 setadd . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48 setremove . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48 shutdown . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66 sin . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 sinh. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 sqrt. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 strcmp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 string hash . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 strsub . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 substitute . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 suspend . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62

T

tan . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 tanh. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 task id. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62 task stack . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 ticks left . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62 time. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 tofloat . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 toint. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38 toliteral . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38 tonum. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38 toobj. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38 tostr. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38 trunc. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 typeof . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37

U

unlisten . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59

V

valid. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 value bytes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 value hash . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 verb args . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 verb code . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 verb info . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 verbs. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52

Table of Contents

1 Introduction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

2 The LambdaMOO Database . . . . . . . . . . . . . . . . . . . . . . . 2

2.1 MOO Value Types . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 2.2 Objects in the MOO Database . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4

2.2.1 Fundamental Object Attributes . . . . . . . . . . . . . . . . . . . . . . . . . 4 2.2.2 Properties on Objects . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 2.2.3 Verbs on Objects . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7

3 The Built-in Command Parser . . . . . . . . . . . . . . . . . . . . . 9

4 The MOO Programming Language . . . . . . . . . . . . . . . 13

4.1 MOO Language Expressions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 4.1.1 Errors While Evaluating Expressions . . . . . . . . . . . . . . . . . . . 13 4.1.2 Writing Values Directly in Verbs . . . . . . . . . . . . . . . . . . . . . . . 14 4.1.3 Naming Values Within a Verb . . . . . . . . . . . . . . . . . . . . . . . . . 14 4.1.4 Arithmetic Operators . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 4.1.5 Comparing Values . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 4.1.6 Values as True and False . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 4.1.7 Indexing into Lists and Strings . . . . . . . . . . . . . . . . . . . . . . . . 19

4.1.7.1 Extracting an Element from a List or String … 19 4.1.7.2 Replacing an Element of a List or String . . . . . . . 20 4.1.7.3 Extracting a Subsequence of a List or String … 21 4.1.7.4 Replacing a Subsequence of a List or String …. 21

4.1.8 Other Operations on Lists . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 4.1.9 Spreading List Elements Among Variables . . . . . . . . . . . . . . 23 4.1.10 Getting and Setting the Values of Properties. . . . . . . . . . . 24 4.1.11 Calling Built-in Functions and Other Verbs . . . . . . . . . . . . 25 4.1.12 Catching Errors in Expressions . . . . . . . . . . . . . . . . . . . . . . . 27 4.1.13 Parentheses and Operator Precedence . . . . . . . . . . . . . . . . . 27

4.2 MOO Language Statements . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 4.2.1 Errors While Executing Statements . . . . . . . . . . . . . . . . . . . . 28 4.2.2 Simple Statements . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 4.2.3 Statements for Testing Conditions . . . . . . . . . . . . . . . . . . . . . 29 4.2.4 Statements for Looping . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 4.2.5 Terminating One or All Iterations of a Loop . . . . . . . . . . . . 32 4.2.6 Returning a Value from a Verb . . . . . . . . . . . . . . . . . . . . . . . . 32 4.2.7 Handling Errors in Statements. . . . . . . . . . . . . . . . . . . . . . . . . 33 4.2.8 Cleaning Up After Errors . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 4.2.9 Executing Statements at a Later Time . . . . . . . . . . . . . . . . . 34

4.3 MOO Tasks. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35 4.4 Built-in Functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36

4.4.1 Object-Oriented Programming. . . . . . . . . . . . . . . . . . . . . . . . . 36 4.4.2 Manipulating MOO Values . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37

4.4.2.1 General Operations Applicable to all Values. … 37 4.4.2.2 Operations on Numbers . . . . . . . . . . . . . . . . . . . . . . 39 4.4.2.3 Operations on Strings . . . . . . . . . . . . . . . . . . . . . . . . 41 4.4.2.4 Operations on Lists . . . . . . . . . . . . . . . . . . . . . . . . . . 47

4.4.3 Manipulating Objects . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48 4.4.3.1 Fundamental Operations on Objects . . . . . . . . . . . 48 4.4.3.2 Object Movement . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 4.4.3.3 Operations on Properties . . . . . . . . . . . . . . . . . . . . . 51 4.4.3.4 Operations on Verbs . . . . . . . . . . . . . . . . . . . . . . . . . 52 4.4.3.5 Operations on Player Objects . . . . . . . . . . . . . . . . . 55

4.4.4 Operations on Network Connections . . . . . . . . . . . . . . . . . . . 55 4.4.5 Operations Involving Times and Dates . . . . . . . . . . . . . . . . . 60 4.4.6 MOO-Code Evaluation and Task Manipulation . . . . . . . . . 60 4.4.7 Administrative Operations . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65

5 Server Commands and Database Assumptions … 67

5.1 Built-in Commands . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 5.1.1 Command-Output Delimiters . . . . . . . . . . . . . . . . . . . . . . . . . . 67 5.1.2 Programming . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 5.1.3 Flushing Unprocessed Input . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 5.1.4 Initial Punctuation in Commands . . . . . . . . . . . . . . . . . . . . . . 69

5.2 Server Assumptions About the Database . . . . . . . . . . . . . . . . . . . . . . . . 69 5.2.1 Server Options Set in the Database . . . . . . . . . . . . . . . . . . . . 69 5.2.2 Server Messages Set in the Database . . . . . . . . . . . . . . . . . . . 70 5.2.3 Checkpointing the Database . . . . . . . . . . . . . . . . . . . . . . . . . . . 71 5.2.4 Accepting and Initiating Network Connections . . . . . . . . . . 71 5.2.5 Associating Network Connections with Players . . . . . . . . . . 72 5.2.6 Out-of-Band Commands . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73 5.2.7 The First Tasks Run By the Server . . . . . . . . . . . . . . . . . . . . 74 5.2.8 Controlling the Execution of Tasks . . . . . . . . . . . . . . . . . . . . . 74 5.2.9 Controlling the Handling of Aborted Tasks . . . . . . . . . . . . . 75 5.2.10 Matching in Command Parsing . . . . . . . . . . . . . . . . . . . . . . . 76 5.2.11 Restricting Access to Built-in Properties and Functions

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76 5.2.12 Creating and Recycling Objects . . . . . . . . . . . . . . . . . . . . . . 76 5.2.13 Object Movement . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77 5.2.14 Temporarily Enabling Obsolete Server Features . . . . . . . . 77

6 Function Index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 78